A Complete Bibliography of Publications in Canadian Journal of Mathematics = Journal canadien de

mathématiques for the decade 1950–1959

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 1.08

Title word cross-reference

3 [349]. [4] [531]. [αn] [194]. a [582]. B [59, 65]. C [453, 46, 109]. C(K) [161]. F [540, 122]. f(f(z)) = g(z) [341]. f(n) [292]. f^{n}(x) = f(x) [156]. F C [200]. B [184]. G [535]. L^{λ} [502]. GF(2) [588, 587]. GL(3, q) [75]. GL(4, q) [75]. L^{1} [332]. L^{2} [503]. L^{p} [226]. λ [502]. λ(τ) [106]. Lu = λMu [507]. m [279]. MT^{∗} [502]. n [68, 489, 292, 25, 513, 29, 192, 164, 194]. p [152]. P G(3, 3) [512, 511]. PGL(3, q) [75]. PGL(4, q) [75]. psbr(n) [406]. q [55]. S [144, 263, 466, 513]. Ssp3 [299]. xspd = 1 [293].

-adic [184]. -chain [200]. -configuration [513]. -conjugates [502]. -convex [46]. -dimensional [25]. -equation [122]. -functions [144, 263, 466]. -local [582]. -measure [453]. -rings [540]. -space [349, 513, 68]. -tensors [152].

A. [327]. abelian [278, 432, 130, 233, 337]. absolute [92]. absolutely [546, 96, 16]. Abstract [430, 40, 7]. abstractly [230]. abstractly-valued [230]. abstrait [7]. accessible [371]. additive [141]. adic [184]. adjoint [228, 508, 250]. admitting [159]. affine [518, 125, 56, 189, 14]. algebra [52, 144, 263, 155, 447]. Algebraic [372, 493, 288, 521, 54, 358, 220, 298]. algebras [295, 235, 59, 65, 561, 198, 102, 569, 325, 24, 469, 245, 362, 537, 568, 564, 448, 196, 120, 313, 539, 143, 535]. algebroid [530]. algorithm [27, 421, 77]. algorithms [495]. almost [362]. alternative [112, 362]. amongst [355]. Analyse [99]. analysis [99]. Analytic [137, 530, 62, 353, 573, 447]. anneau [7]. annulus [136]. application [421, 306]. applications [96, 574, 161, 55]. approximation [560, 23, 26, 431, 493]. Approximations [25, 578]. arbitrary [355]. arc [124]. arcs [349]. area [426, 25]. arithmetic [153, 38, 8]. associated [166, 531, 30]. asymptoic [360]. Asymptotic [286, 348, 359, 420, 410, 358, 328, 579]. atmosphere [148, 303]. Aubry [90]. Autometrization [181]. Autometrized [59, 65]. automorphes [436, 497]. Automorphic [436, 497]. automorphisms [40, 326]. averages [58]. averaging [481, 171, 385]. axiom [393]. Axiomatique [50]. Axioms [107].

Babinet [529]. Baer [364]. balanced [214, 189, 241, 312]. Banach [561, 245, 355, 448]. barrier [20]. base [110]. basis [381]. be [134]. behavior [328]. behaviour [286, 579]. Bernoulli [418, 367]. Bernstein [207, 158]. Bessel [249, 57, 549, 157, 548]. Betti [499]. between [178, 21, 260, 516, 485, 120]. Bianchi [12]. bilateral [205, 487]. binary [257, 543]. binomial [440]. bipartite [517]. block [214, 189, 241, 312]. bodies [556]. Boolean [404, 52, 197, 59, 65, 569, 155, 196, 164]. Borel [391]. boundaries [128]. Boundary [166, 228, 252, 357, 347, 508, 383, 251, 250]. bounded [527, 409, 583, 524]. bounding [150]. bounds [426, 282, 382]. Brauer [327]. Brouwerian [232]. Burkhardt [531]. Burkill [480]. Burnside [437].

calculus [123]. can [134]. canonical [118, 279, 333]. canonique [333]. capacity [272]. caractérisation [332]. Cartan [351]. Cartesian [539]. Cartwright [265]. categories [461]. Cauchy [412, 383, 268]. Cayley [460, 534]. Central [491]. Certain [380, 64, 229, 439, 336, 446, 487, 541, 395, 494, 91, 102, 21, 131, 220, 79, 259, 355, 360, 170, 338, 38, 88, 30, 189]. certains [226, 88]. chain [200]. chains [115, 544]. changes [91]. channel [20]. character [468, 244]. Characteristic [405, 329]. characterization [197, 332]. characterizations [398, 128]. characterizes [456]. Characters [539, 467, 183, 264]. chords [512, 511]. chromatic [273, 219]. Chung [132]. circulants [407]. class [97, 212, 257, 276, 387, 477, 533, 562, 300, 442, 294, 362, 360, 590, 18, 313, 544, 337]. Classes [407, 56, 439]. classiques [238]. closed [447]. closure [258, 117, 165]. coclosed [287]. codes [488]. coefficient [440, 136]. coefficients [459, 523, 106, 579]. cognate [205]. cohomology [198, 401]. collineations [298]. colour [142, 424]. combinations [207]. Combinatorial [10, 273, 441, 574, 361]. Comma [488]. Comma-free [488]. commutative [73]. commutativity [316, 463, 170]. commuting [365]. compact [161, 62, 450]. compactes [161]. compactness [304, 56]. complementary [455]. Complemented [576]. Completeness [215, 462, 444]. complex [333]. Complexes [233, 333]. concernant [577]. Concerning [543, 190, 554, 297, 461, 583, 577, 165]. condition [131, 463, 344]. conditions [384, 433, 146, 383]. cones [398]. configuration [83, 531, 513, 100]. configurations [100]. Conformal [305, 201, 349]. Congruence [521, 192]. Congruences [459, 311, 521, 523, 202, 17]. Conics [374]. conjecture [154, 528, 132]. conjugates [502, 501, 131]. connected [418, 86, 543]. connectedness [377]. connection [120]. connections [351]. connectivity [557]. consecutive [37]. considerations [323]. consistency [213]. Constant [151, 149]. constraints [89]. construction [584, 224]. Constructions [376]. containing [348]. continua [48]. Continued [546, 100]. continues [100]. Continuous [156, 540, 4, 580, 251]. Contribution [36, 219, 36]. Contributions [103]. convergence [503, 227, 478]. convergent [546, 96]. Convex [375, 580, 79, 56, 577, 23, 46, 556]. convexes [577]. Convolution [284, 159]. coordinate [146]. coordinatization [319]. Core [213]. Core-consistency [213]. Correction [425, 591, 174, 515, 289, 254]. Corresponding [149]. Countable [330]. countably [74, 449]. courbure [29, 116]. covering [366]. Coverings [517, 167, 301]. criterion [172, 117, 165]. cubes [475]. cubic [588, 587, 586]. curvature [151, 177, 29, 116, 231, 289, 149]. curve [231, 289]. curves [68, 375, 221, 220, 202, 44, 493, 530]. Cyclic [98, 125, 212, 538, 77, 67, 220, 298, 436]. Cyclically [445]. cyclique [436]. cyclotomic [97, 255].

D. [144]. dans [99]. Darboux [96]. data [268]. decimals [104]. decompositions [34]. Dedekind [326, 370, 444]. defined [274, 176]. definite [427, 407]. definitions [430]. degree [121, 276, 302]. degrees [16]. Denjoy [528]. Dense [557]. denumerable [110]. depth [20]. derivatives [271, 354, 505]. Desargues [318]. Desarguesian [442, 393]. designs [214, 189, 241, 312]. d’espaces [161]. determinants [361]. determined [209, 489]. developments [248]. diagonable [372]. diagonal [102]. diagram [262]. diagrams [9]. difference [554, 181, 191, 119, 190, 473]. Differentiable [201, 349]. Differential [267, 270, 390, 95, 507, 425, 414, 87, 41, 218, 415, 345, 160, 329, 348, 508, 485, 383, 88, 30, 250, 134, 159, 204, 254]. differentiation [300]. différentiels [88]. Dimension [496, 403, 456]. dimensional [158, 25]. dimensions [29]. diophantine [308, 22, 336, 26, 431, 422]. Direct [76]. Dirichlet [483]. disconjugate [345]. Discrete [82]. discriminants [257]. Disjoint [559]. disorientation [475]. dissimilar [516]. Distance [71, 59]. distance-theoretic [59]. distinct [492]. distortion [305]. distributed [522]. distribution [206, 281, 35, 310, 99]. distributions [64, 229, 99, 344]. distributive [51]. distributives [50]. divergent [139]. divisible [344]. division [495, 501, 81, 170, 143, 535]. divisors [256]. Doetsch [506]. domain [348]. domains [23]. Double [394, 253, 503, 272]. dual [569, 18]. dualities [318]. duality [68]. d’un [7, 29, 520]. duodenary [187]. dynamical [302]. dynamics [13, 49].

each [257]. effectif [520]. Einstein [177]. elastic [384]. element [313]. Elementarfunktionen [15]. elementary [243, 568, 8, 15]. elements [404, 108, 105, 524, 296, 339]. eleven [67]. ellipsoid [25]. elliptic [203, 357, 108, 107, 513, 552]. embedding [214]. embeddings [266]. endo [497]. endomorphisms [365]. ennuple [202]. ensemble [520]. Entire [341]. entre [238]. Enumeration [379]. Enveloping [24]. equation [356, 507, 122, 166, 203, 357, 347, 156, 21, 160, 268, 134, 341, 16, 43, 43]. equations [270, 267, 390, 87, 408, 269, 41, 218, 415, 482, 553, 129, 160, 22, 72, 146, 329, 348, 485, 145, 383, 336, 88, 308, 30, 250, 204, 254, 145]. equivalence [56, 326, 530, 458]. equivalence-classes [56]. Equivalent [432]. Errata [61]. error [91, 281]. espace [29]. espaces [226, 332, 333, 99]. étude [36]. Euclid [27, 77]. Euclidean [108, 513]. Euler [281]. even [546]. example [520]. exceptional [259]. exceptionnels [238]. Exemple [520]. existence [281, 480, 110, 344, 189]. expansion [269]. Expansions [359, 420]. explicit [282, 382]. exponential [54, 367]. exponentiels [367]. expressions [425, 414]. Extension [167, 377, 327, 111, 174, 354, 391, 534]. Extensions [198, 541, 259, 334]. Extremal [392, 567, 124]. Extreme [93, 291, 428, 187, 290]. Extremum [3].

facially [31]. facially-regular [31]. factor [212, 242]. Factorization [465, 6]. factorized [134]. factors [550, 307, 127]. faiblement [161]. family [141, 110, 473]. Fan [435]. Farey [55]. feasibility [574]. field [97, 180, 212, 395, 408, 49, 129, 54, 329, 2, 42]. fields [288, 27, 521, 495, 77, 297]. finding [421]. finis [238]. Finite [126, 60, 237, 180, 395, 374, 129, 274, 20, 199, 184, 326, 301, 536, 394, 317, 396, 438, 6, 242, 16, 143]. Finsler [29, 389]. finslérien [29]. first [212, 269, 482, 231, 289]. five [220]. fixed [265, 299, 444]. fixed-point [444]. Flächen [15]. flat [518]. flow [378, 474, 574]. flows [421]. Focal [199]. fonctions [577]. form [113, 187, 427, 49, 333]. formal [240]. forme [333]. forms [291, 428, 243, 257, 93, 287, 111, 174, 397, 290, 62, 459, 523, 130, 134, 458]. formula [468, 285, 342, 159, 150]. formulations [52]. Fourier [321, 122, 206, 503, 64, 229, 551, 106, 284]. fourth [250]. fractional [39]. fractions [546, 100]. Fréchet [34]. free [494, 538, 488, 363]. freedom [302]. French [63, 498, 563, 161, 332, 333, 7, 577, 145, 29, 116, 88, 520, 436, 497, 99, 36, 100, 367, 43]. frequencies [4, 5]. Fresnel [455, 413]. Frobenius [464, 182, 533]. Full [300]. Function [209, 175, 541, 228, 154, 381, 571, 91, 281, 459, 106, 211]. functional [15, 156, 452, 343, 341]. functionals [162, 208]. Functions [505, 15, 163, 417, 540, 545, 271, 144, 263, 206, 426, 137, 283, 21, 117, 141, 527, 80, 354, 583, 62, 455, 4, 5, 466, 140, 434, 568, 577, 101, 38, 446, 157, 353, 136, 230, 551, 573, 110, 447, 30, 164, 504, 188, 548, 547]. Fundamental [59]. Funktionentheorie [15]. Further [523, 534].

G. [577]. Galois [408, 73, 120]. Games [450, 380]. gamma [211]. Gauss [451, 526]. General [205, 553, 129, 21]. généralisation [116, 29]. generalization [83, 262, 41, 412, 116, 231, 289, 352, 29]. generalizations [492, 437]. generalize [160]. Generalized [295, 13, 527, 481, 12, 555, 571, 542, 28, 409, 505, 479, 177, 564, 484, 385, 128]. generate [494]. generated [429, 186]. Generating [157, 548, 547, 417, 154, 274]. genus [257, 66, 168, 575]. Geodesic [476, 218]. geometric [497]. geometries [126, 393, 11, 14]. géométrique [497]. geometry [95, 514, 108, 107, 376, 94, 232]. German [15]. girth [275]. given [89, 40, 516, 457]. Glaisher [179]. global [399]. Goursat [470]. Graph [532, 275, 121, 516, 457]. graph-subgraph [516]. graph-theoretical [457]. Graphs [457, 517, 239, 379, 273, 516, 6, 127, 242, 586]. Graves [165]. gravitation [12, 177, 145]. gravitational [72, 145]. Green [228, 62, 30]. Gross [222]. group [85, 588, 587, 65, 239, 40, 198, 200, 294, 53, 105, 133, 183, 264, 261, 314, 457, 436, 497, 9, 16]. groupe [436, 497]. groupes [238, 498, 563]. Groupoïdes [436, 497]. groupoids [436, 497]. Groups [429, 388, 131, 277, 395, 494, 86, 499, 538, 498, 563, 278, 464, 274, 276, 533, 562, 476, 199, 240, 294, 467, 184, 326, 293, 185, 536, 400, 446, 431, 233, 186, 237, 430, 69, 396, 438, 337, 589, 363, 432].

H [132]. Hadamard [384]. half [573, 589]. half-groups [589]. half-plane [573]. halo [300]. Hamiltonian [13, 49]. Hardy [353, 352]. Harmonic [152, 353, 287, 416, 393, 99]. harmonics [225]. harmoniques [99]. Hasse [472]. Hausdorff [109]. having [218]. Hayman [583]. heat [160]. Helly [23]. Hermite [417, 547]. Hermitian [433, 333, 392, 434, 567]. hermitienne [333]. hexagonal [309]. Higman [364]. Hilfsmittel [15]. Hill [43]. Hitchcock [421]. holomorphic [151]. homogeneous [333]. homogènes [333]. homology [278]. homomorphic [235, 102]. hook [239]. horocyclic [431]. hulls [375]. hyper [399]. hyper-surfaces [399]. Hyperbolic [416, 514, 152, 347, 415, 553, 376, 268]. hyperconvexity [581]. hyperelliptic [118]. hypergeometric [205, 487, 525]. hypersurface [231, 289, 44]. Hypersurfaces [389].

ideal [279, 2, 103]. ideals [496, 569, 439, 447]. Identities [17, 12, 523]. II [521, 65, 477, 283, 28, 76, 567, 303, 420, 264, 342, 526, 438, 45]. III [498]. images [235]. imaginary [495]. Imbedding [433]. immersibility [53]. immersion [565]. Implications [384]. implicative [197]. improved [554]. Incidence [45, 98]. inclusion [213]. incomplete [214, 189, 241, 312]. indefinite [111, 174, 397]. independent [35]. index [124]. individual [300]. Induced [324, 33]. Inequalities [434, 246, 247]. inequality [352]. Infeld [134]. infinite [380, 294, 307, 544]. infinitely [344]. inhomogeneous [309]. inner [466]. integer [267, 454]. integers [550, 234, 292, 37, 171, 194]. integrable [271, 453]. Integral [163, 258, 253, 21, 411, 28, 480, 184, 454, 230]. integrals [96, 452, 527, 455, 413, 552, 150]. integration [504, 188]. integro [159]. integro-differential [159]. interchangeable [491]. intersection [79]. interval [153]. invariance [568]. invariant [220, 298, 472]. invariants [60, 33]. inverse [286]. inversion [386, 138, 230, 285, 342, 451, 526, 159]. investigations [361]. involution [67, 448]. Involutions [531, 220, 299]. irrationality [172]. irreducibility [323]. irreducibilty [556]. irreducible [133, 16]. irregular [137]. isomorphic [363]. isomorphismes [238]. Isomorphisms [307, 245]. Isoperimetric [123]. Iterated [114]. Iterates [39]. iteration [38]. IV [563, 261].

J. [319, 132]. Jacobi [178, 41]. Jordan [235].

kind [552]. kinematics [108]. knots [1]. known [426]. Köthe [208]. Kronecker [472]. Kummer [509]. Ky [435].

labelled [379]. Laguerre [417, 484].

Landau [382]. Laplace

[166, 286, 138, 230, 285, 342, 484]. Lattice [542, 266, 195, 519]. Lattices [40, 576, 173, 236, 50, 560, 478, 51, 120, 462]. Laurent [579]. Laves [275]. least [305]. Lebesgue [63]. left [34]. Legendre [322]. Lehmer [486]. lemma [288, 470]. length [209]. levelling [209]. Lie [499, 468, 498, 563, 198, 24]. limit [491]. limits [409, 34]. line [94, 268]. linéaires [161]. lineal [108]. Linear [207, 537, 568, 246, 167, 270, 390, 356, 311, 395, 87, 408, 521, 41, 218, 357, 415, 482, 89, 452, 161, 82, 208, 328, 383, 336, 192, 247, 14, 396, 438, 120, 566, 256]. linear-space-valued [167]. linearly [137]. lines [489]. Liouville [454]. Littlewood [144, 265, 353, 352]. Local [377, 582]. locally [301, 453, 6]. Logarithmic [272]. logics [369]. Lorentz [306]. lower [426]. manifold [452]. manifolds [554, 581, 128]. Map [142, 424]. Map-colour [142, 424]. mapping [102, 453]. mappings [167]. Maps [220, 305, 575, 18]. Maschke [370]. Mathieu [400, 69]. matric [102]. Matrices [404, 494, 433, 279, 102, 98, 392, 434, 567, 537, 568, 524, 441, 472, 296, 339]. matrix [295, 228, 558, 327, 340, 591, 182, 471]. Maurice [570]. Maximal [378, 361, 421, 353, 256]. maximum [425, 414, 35, 435]. McCoy [182]. mean [478]. measurable [137]. Measure [135, 391, 453]. measures [304, 584]. medians [514]. mehrerer [15]. ménages [510]. meromorphic [62]. Metamathematical [323]. method [246, 391, 247, 134]. methods [340, 591, 176, 76, 213]. metric [555, 332, 456]. métrique [332]. Metrization [70]. Meyer [174, 111]. Minimal [22, 476]. minimum [309, 425, 414, 427]. minimum-maximum [425, 414]. Minkowski [472]. Mixed [482, 553, 252, 416, 415]. Modified [357]. Modular [85, 576, 185, 338, 459, 523, 261, 314, 106]. modules [324, 326, 370]. modulo [279]. moments [344]. monotone [577]. monotones [577]. monotonic [354]. Morse [571]. motion [72, 146]. motions [65]. multicoherent [45]. multigroups [11]. Multiple [147, 15]. multiplicative [162]. Multiply [140]. mutables [15].

Nagao [223]. nature [87]. nearly [487]. nearly-poised [487]. negative [163, 524]. nets [60]. Network [474, 378, 421, 574]. network-flow [574]. Neumann [319]. Nikodym [63]. nilpotent [498, 563, 294]. nilpotents [498, 563]. Noetherian [402]. Nombres [367]. Non [216, 393, 73, 566, 270, 267, 512, 96, 442, 508, 171, 524, 520, 189, 511]. non-absolutely [96]. non-averaging [171]. non-commutative [73]. Non-Desarguesian [393, 442]. non-existence [189]. non-integer [267]. Non-linear [566, 270]. non-negative [524]. Non-normal [73]. Non-parametric [216]. non-projectif [520]. non-projective [520]. non-ruled [512, 511]. non-self-adjoint [508]. non-semisimple [73]. noncommutative [103]. nonconvex [309]. nonharmonic [284]. nonnormal [387, 477]. Normal [89, 115, 104, 152, 415, 433, 449, 73]. normality [260]. normalizer [338]. norms [322]. Note [291, 104, 263, 327, 192, 185, 386, 564, 30, 32, 211, 331, 410, 495, 581, 400, 139, 241, 84]. nouvelles [145]. Null [204, 254, 585]. number [97, 212, 521, 495, 195, 129, 516, 54, 489, 590, 8, 164, 490]. numbers [418, 499, 260, 255, 367]. Numerical [188, 60].

occurring [544]. odd [546]. Old [1]. omitted [283]. one [121, 336, 302]. one-regular [121]. ones [441]. onto [102]. open [155]. operations [52]. operator [508, 385, 582]. operators [481, 388, 387, 477, 544]. order [267, 390, 87, 269, 68, 218, 347, 415, 482, 553, 215, 39, 255, 268, 454, 590, 250, 544, 16]. ordered [445, 331, 266, 82, 543]. orders [325]. ordinary [205, 489, 485, 250]. Oriented [518]. origin [2]. Orthogonal [363, 525, 374, 117, 467, 202]. orthogonality [536]. orthonormal [446]. oscillatoires [145]. oscillatory [145]. Osima [223]. Ovals [318, 317].

pair [269, 516]. pairs [494, 408]. paper [591, 570, 32]. Pappus [83]. parabolic [268]. paracompact [74, 449]. paracompactness [330]. Parallel [221]. parallelogram [195]. parameterization [460]. parametric [216, 308]. parametrization [534]. part [546, 261]. partial [218, 415, 160, 329, 35, 485, 383]. partially [266, 312]. partition [459]. Partitioning [153]. partitions [180, 542]. Pencils [350, 585]. perfect [428, 84]. period [429, 67]. Periodic [14, 88]. périodiques [88]. periods [220]. permutation [277, 276, 360]. Perron [411]. person [380]. perturbation [270]. perturbed [88]. perturbés [88]. Petrie [490]. phenomenon [259]. plane [374, 20, 201, 148, 303, 393, 573, 317]. planes [125, 442, 394, 443, 515]. plethysm [466]. plethysms [467]. Point [301, 265, 348, 393, 250, 444]. Point-finite [301]. Points [210, 195, 137, 489, 201, 349, 299]. poised [487]. polarities [350, 585]. polygon [490]. polyhedra [31, 78, 149, 3]. polyhedral [398]. polynomes [7]. polynomial [373, 496, 178, 7, 439, 334]. polynomials [249, 417, 419, 57, 207, 158, 418, 554, 323, 322, 549, 157, 484, 405, 219, 368, 525]. Polytopes [588, 224]. positive [388, 154, 407]. possible [134]. postage [36]. postage-stamp [36]. poste [36]. Postulates [51, 196]. potential [287]. power [358, 240, 55, 396, 438, 244]. Practical [520]. Prime [569, 396, 438, 550, 276, 329, 292, 8, 256, 194]. prime-number [8]. primes [243]. primitive [195]. principal [412, 279, 29, 116]. principle [529, 435]. principles [584]. probability [532]. problem [425, 414, 252, 415, 381, 421, 253, 572, 354, 508, 492, 383, 268, 23, 422, 136, 109, 330, 250, 36, 100, 582, 43]. problème [36, 100, 510, 43]. problems [10, 228, 166, 357, 347, 416, 482, 553, 123, 360, 251]. processes [491]. produced [324]. product [509, 472]. Products [467, 453, 538, 561, 105, 549, 539]. progressions [8]. projectif [520]. projection [452]. projective [126, 442, 393, 394, 443, 515, 520, 350, 317, 14]. projectivities [374]. Proof [154, 113, 243, 424, 265, 486, 8, 242]. propagation [148, 303]. properties [59, 96, 300, 358, 79, 392, 567, 573, 441, 457, 3, 46]. property [55, 109]. Pseudo [315]. Pseudo-regularity [315]. pseudometrics [167].

quadratic [243, 180, 27, 257, 408, 521, 427, 495, 322, 111, 174, 397, 259, 130, 458]. quadric [512, 511]. Quadrics [587]. quantization [147]. quasi [356, 357, 583]. quasi-bounded [583]. quasi-linear [356, 357]. Quasiconvex [47]. quasigroupes [497]. quasigroups [497]. quaternion [296, 339]. Quelques [116]. quiet [148, 303].

Radiation [72]. Radicals [373].

Ramanujan [17]. random [475, 35]. rank [471]. ranks [558]. real [545, 27, 296, 339]. recurrences [256]. recursions [86]. recursive [566]. Reducible [308]. reflection [237, 430]. reflections [429, 186]. regions [309, 283, 56]. Regular [168, 31, 78, 66, 121, 340, 591, 348, 343, 155, 136, 575, 3]. regularity [315]. related [281, 157, 284]. relation [178, 21]. relations [274, 273, 260, 300, 485, 543, 536, 45]. relationships [355]. Relative [401, 323]. relatively [292, 194]. relaxation [246, 247]. relevance [588, 587]. remark [536, 110, 165]. Remarks [274, 550, 316, 461, 290, 116, 183, 264, 402]. remarques [116]. Representation [208, 448, 243, 469, 536, 314, 138, 230, 551]. representations [85, 521, 498, 563, 37, 184, 326, 185, 133, 590, 33, 261, 308, 75, 396, 438, 16, 363, 498, 563]. representatives [474, 492]. residue [191, 244]. residues [55, 440]. resolvable [189]. Resolvents [395]. respect [384, 527]. restricted [383]. result [554]. results [578, 328, 1]. Retraceable [500]. Riemann [113, 15, 271, 176, 505, 454]. Riemannian [289, 320, 152, 581, 202, 231]. Riemannschen [15]. Ring [245, 404, 52, 279, 294, 7, 369]. ring-logics [369]. rings [112, 373, 496, 541, 540, 197, 365, 372, 501, 102, 169, 316, 463, 81, 371, 326, 465, 439, 73, 170, 370, 334, 335, 402]. Robinson [32]. Roch [113]. rotation [538]. Roux [346]. ruled [512, 511].

samples [89]. satisfy [393]. satisfying [21]. scalaire [29, 116]. scalar [29, 116]. Schlicht [483]. Schoenberg [154]. Schottky [282]. Schrödinger [134]. second [390, 87, 218, 347, 415, 268]. second-order [390]. sector [34]. Selberg [288]. selections [580]. self [228, 508, 18, 250]. self-adjoint [228, 250]. self-dual [18]. semi [82, 24, 4, 462]. semi-continuous [4]. semi-lattices [462]. semi-ordered [82]. semi-simple [24]. semicontinuous [5]. semigroup [53]. semigroups [135]. semisimple [73]. senary [428]. separability [331]. separable [325]. Separation [560]. sequence [154]. sequences [409, 522, 355, 55, 566]. Series [549, 205, 410, 503, 509, 199, 358, 240, 248, 360, 139, 423, 483, 272, 487, 312, 284, 204, 254, 525]. Set [277, 279, 137, 409, 450, 520]. Set-transitive [277]. Sets [423, 445, 266, 178, 500, 47, 71, 234, 79, 191, 105, 119, 171, 190, 155, 272, 473, 46]. seven [220]. several [353, 188]. shock [148, 303]. Short [52, 424, 265, 242]. sides [426, 490]. Siegel [193]. sign [91]. simple [112, 499, 375, 421, 24, 150]. Simultaneous [408, 521, 26]. single [257]. singular [87, 228, 554, 269, 137, 348, 508, 343, 251]. Sion [570]. sixteen [255]. sixteenth [244]. solution [348]. solutions [390, 218, 156, 129, 22, 145, 293, 88, 341]. solvable [562]. Some [162, 417, 578, 418, 550, 260, 584, 399, 358, 461, 92, 119, 328, 492, 406, 183, 264, 312, 46, 128, 296, 437, 402, 535, 129, 182, 485, 248, 459, 306, 506, 573, 84, 368, 116]. sous [226]. sous-espaces [226]. Souslin [330]. space [167, 331, 366, 68, 375, 555, 518, 94, 349, 513, 355, 29, 202, 231, 289, 389, 350, 99, 120]. Spaces [403, 377, 70, 320, 210, 74, 152, 209, 381, 502, 571, 161, 332, 82, 175, 333, 543, 208, 449, 355, 149, 306, 45]. special [235, 468, 487]. spectra [251]. Spectral [507, 387, 477, 228]. spectrum [87]. sphere [108, 101]. spheres [366]. Spherical [11, 225]. Spinor [94]. square [19]. squared [84]. squares [84]. Squaring [19]. stability [384]. Stable [173, 236]. stamp [36]. Standard [371]. Star [9, 309]. star-regions [309]. statistics [215]. Steinitz [222]. step [446]. Stieltjes [321, 411, 386]. stochastic [558]. Störmer [43]. structure [464, 589]. structures [580]. study [36]. sub [423]. sub-series [423]. subadditive [140]. subalgebras [469, 143]. subgraph [516]. subgraphs [557]. subgroups [338]. submanifolds [518]. submerged [20]. Submethods [340, 591]. subsets [559]. Subspaces [320, 555, 461, 120]. Substitution [240]. successive [578]. sum [380, 48, 37]. Summability [176, 340, 591, 76, 213, 92]. sums [222, 176, 54, 35, 234]. Supersoluble [565]. supremum [141]. surface [588, 587, 575]. surfaces [15, 66, 118, 168, 476, 399, 298]. symmetric [113, 86, 85, 181, 239, 426, 467, 434, 568, 192, 293, 185, 133, 183, 264, 261, 314, 9, 339]. symmetries [225]. symmetry [309, 164, 586]. symplectic [467]. synthèse [99]. synthesis [99]. system [118]. systèmes [88]. Systems [311, 160, 41, 482, 474, 345, 328, 302]. Szekeres [577].

Tauberian [479, 454]. Tchebycheff [23]. ten [275]. Tensor [561, 166, 203, 252, 177]. tensors [152, 416]. term [558, 281, 471]. termwise [298]. ternary [427, 111, 174, 397]. th [55]. their [588, 587, 300, 308]. theorem [193, 113, 222, 90, 243, 179, 269, 41, 68, 346, 63, 424, 223, 319, 574, 327, 280, 503, 214, 265, 169, 297, 282, 382, 111, 174, 81, 583, 479, 470, 56, 577, 299, 318, 454, 353, 486, 8, 364, 242, 437, 444, 335]. théorème [63, 7, 577]. theorems [491, 142, 384, 182, 300, 399, 76, 92, 119, 406, 23, 506, 296]. theoretic [59]. theoretical [457]. theory [115, 15, 507, 287, 278, 12, 532, 216, 387, 477, 542, 80, 177, 7, 248, 227, 103, 536, 73, 138, 230, 544, 219, 42, 369]. third [198, 552]. Three [31, 66, 121, 297, 220, 149, 575]. Thue [90]. timbres [36]. tool [15]. topological [70, 80, 560]. topologies [519]. total [213]. totally [154]. totatives [310]. Touchard [419, 418, 368]. transfer [7]. transfert [7]. transfiniment [520]. transfinitely [520]. transform [286, 159]. transformation [138, 285, 342, 451, 526]. transformations [205, 537, 568, 217, 487, 14, 120]. transforms [321, 122, 206, 114, 21, 64, 229, 484, 551, 284]. transitive [277]. Transitivities [443, 515]. transitivity [394]. Transue [571]. transversals [559]. triangle [514, 426]. trigonometric [272, 204, 254]. trigonometrical [205]. Two [78, 316, 290, 158, 429, 168, 380, 426, 475, 509, 250]. two-dimensional [158]. two-person [380]. two-point [250]. type [205, 203, 252, 161, 469, 572, 307, 386, 17, 487, 159]. types [129, 260, 164]. Typically [545]. Typically-real [545].

ultraspherical [178]. underlying [584]. Unified [42]. unimodular [407]. Union [44]. Uniqueness [347, 578, 384, 30]. unit [113]. unitaires [563]. unitaries [498]. Unitary [217, 186, 429, 498, 563, 101, 237, 339]. unity [313]. univalent [283]. unknown [336]. Upper [426]. using [52].

V [63, 314]. value

[228, 166, 252, 357, 347, 412, 508, 251]. valued [167, 381, 230, 504]. values [478]. variables [35, 192, 353, 164, 188]. variances [89]. Variation [202, 527, 34]. variational [253]. variations [123]. varieties [62]. various [260]. vector [304, 381, 48, 560, 504, 120]. vector-valued [504]. vectorial [332]. vectorial-metric [332]. vectorielle [332]. vectorielle-métrique [332]. vectoriels [226]. vectors [222]. Veränderlichen [15]. vertex [78]. vertex-regular [78]. viewpoint [95]. Vitali [300]. Vorticity [58].

Ward [411]. Waring [193]. Water [20]. Watson [21]. waves [20]. Weak [304]. weakly [161]. weakly-compact [161]. Wedderburn [280]. Weighted [180, 322]. Well [522]. which [494, 160, 505, 393, 134]. whose [118, 426]. without [313]. Witt [564]. Wythoffian [224].

Young [262].

Zariski [335]. Zassenhaus [470]. zero [403, 380]. zero-sum [380]. zeros [390, 163, 413, 441]. Zeta [101].

References

Seifert:1950:ONR

[1] Herbert Seifert and William Threlfall. Old and new results on knots. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2(??):1–15, ???? 1950. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mann:1950:FOI

[2] H. B. Mann. On the field of origin of an ideal. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2(??):16–21, ???? 1950. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Toth:1950:EPR

[3] László Fejes Tóth. Extremum properties of the regular polyhedra. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2(??): 22–31, ???? 1950. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Levi:1950:FSC

[4] F. W. Levi. On frequencies and semicontinuous functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2(??):32–43, 1950. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Levi:1950:FSF

[5] F. W. Levi. On frequencies and semicontinuous functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2(??):32–43, ???? 1950. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Tutte:1950:FLF

[6] W. T. Tutte. The factorization of locally finite graphs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2(??):44–49, ???? 1950. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lesieur:1950:TTA

[7] Léonce Lesieur. Un théorème de transfert d’un anneau abstrait à l’anneau des polynomes. (French) [A theory of the transfer of an abstract ring to a polynomial ring]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2(??):50–65, ???? 1950. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Selberg:1950:EPP

[8] Atle Selberg. An elementary proof of the prime-number theorem for arithmetic progressions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2(??):66–78, ???? 1950. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Staal:1950:SDS

[9] R. A. Staal. Star diagrams and the symmetric group. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2(??):79–92, ???? 1950. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chowla:1950:CP

[10] S. Chowla and H. J. Ryser. Combinatorial problems. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2(??):93–99, ???? 1950. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Prenowitz:1950:SGM

[11] Walter Prenowitz. Spherical geometries and multigroups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2(??):100–119, ???? 1950. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Einstein:1950:BIG

[12] A. Einstein. The Bianchi identities in the generalized theory of gravitation. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2 (??):120–128, ???? 1950. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dirac:1950:GHD

[13] P. A. M. Dirac. Generalized Hamiltonian dynamics. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2(??):129–148, ???? 1950. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Snapper:1950:PLT

[14] Ernst Snapper. Periodic linear transformations of affine and projective geometries. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2(??):149–151, ???? 1950. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Behnke:1950:ERF

[15] Heinrich Behnke and Karl Stein. Elementarfunktionen auf Riemannschen Flächen als Hilfsmittel für die Funktionentheorie mehrerer Veränderlichen. (German) [Elementary functions on Riemann surfaces as a tool for the functional theory of multiple mutables]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2 (??):152–165, ???? 1950. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Zassenhaus:1950:EDA

[16] Hans Zassenhaus. An equation for the degrees of the absolutely irreducible representations of a group of finite order. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2(??):166–167, ???? 1950. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ramanathan:1950:ICR

[17] K. G. Ramanathan. Identities and congruences of the Ramanujan type. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2(??): 168–178, ???? 1950. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Smith:1950:CSD

[18] C. A. B. Smith and W. T. Tutte. A class of self-dual maps. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2(??):179–196, ???? 1950. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Tutte:1950:SS

[19] W. T. Tutte. Squaring the square. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2(??): 197–209, ???? 1950. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Heins:1950:WWC

[20] Albert E. Heins. Water waves over a channel of finite depth with a submerged plane barrier. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2(??):210–222, ???? 1950. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Goodspeed:1950:RBF

[21] F. M. Goodspeed. The relation between functions satisfying a certain integral equation and general Watson transforms. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2(??):223–237, ???? 1950. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Holzer:1950:MSD

[22] L. Holzer. Minimal solutions of Diophantine equations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2(??):238–244, ???? 1950. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rademacher:1950:HTC

[23] Hans Rademacher and I. J. Schoenberg. Helly’s theorems on convex domains and Tchebycheff’s approximation problem. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2 (??):245–256, ???? 1950. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Jacobson:1950:EAS

[24] N. Jacobson. Enveloping algebras of semi-simple Lie algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2(??): 257–266, ???? 1950. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lehmer:1950:AAD

[25] D. H. Lehmer. Approximations to the area of an n-dimensional ellipsoid. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2 (??):267–282, ???? 1950. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Raisbeck:1950:SDA

[26] Gordon Raisbeck. Simultaneous Diophantine approximation. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2(??):283–288, ???? 1950. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chatland:1950:EAR

[27] H. Chatland and H. Davenport. Euclid’s algorithm in real quadratic fields. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2(??): 289–296, ???? 1950. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

James:1950:GII

[28] R. D. James. A generalized integral. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2(??):297–306, ???? 1950. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Moor:1950:GDS

[29] Arthur Moór. Généralisation du scalaire de courbure et du scalaire principal d’un espace finslérien à n dimensions. (French) [Generalization of the scalar curvature and the principal scalar of a Finsler space in n dimensions]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2(??): 307–313, ???? 1950. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sears:1950:NUG

[30] D. B. Sears. Note on the uniqueness of the Green’s functions associated with certain differential equations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2(??): 314–325, ???? 1950. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

ApSimon:1950:TFR

[31] Hugh ApSimon. Three faciallyregular polyhedra. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2(??):326–330 (2 plates), ???? 1950. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Todd:1950:NPR

[32] J. A. Todd. Note on a paper by Robinson. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2(??):331–333, ???? 1950. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Robinson:1950:IRI

[33] G. de B. Robinson. Induced representations and invariants. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2(??):334–343, ???? 1950. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Morse:1950:FVS

[34] Marston Morse and William Transue. The Fréchet variation, sector limits, and left decompositions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2(??):344–374, ???? 1950. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kac:1950:DMP

[35] Mark Kac and Harry Pollard. The distribution of the maximum of partial sums of independent random variables. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2(??): 375–384, ???? 1950. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Touchard:1950:CED

[36] Jacques Touchard. Contribution à l’étude du problème des timbres poste. (French) [Contribution to the study of the postage-stamp problem]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2(??): 385–398, ???? 1950. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

LeVeque:1950:RSC

[37] W. J. LeVeque. On representations as a sum of consecutive integers. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2(??): 399–405, ???? 1950. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Niven:1950:ICA

[38] Ivan Niven. The iteration of certain arithmetic functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2(??):406–408, ???? 1950. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Isaacs:1950:IFO

[39] Rufus Isaacs. Iterates of fractional order. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2(??):409–416, ???? 1950. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Frucht:1950:LGA

[40] Robert Frucht. Lattices with a given abstract group of automorphisms. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2(??): 417–419, ???? 1950. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cramlet:1950:GTJ

[41] Clyde M. Cramlet. A generalization of a theorem of Jacobi on systems of linear differential equations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2(??):420–426, ???? 1950. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wyman:1950:UFT

[42] Max Wyman. Unified field theory. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2 (??):427–439, ???? 1950. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

deVogelaere:1950:EHP

[43] René de Vogelaere. Équation de Hill et problème de Störmer. (French) [Hill’s equation and Störmer’s problem]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2(??): 440–456, ???? 1950. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Springer:1950:UCH

[44] C. E. Springer. Union curves of a hypersurface. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2(??):457–460, ???? 1950. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Stone:1950:IRM

[45] A. H. Stone. Incidence relations in multicoherent spaces. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2(??):461–480, ???? 1950. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Valentine:1950:SPC

[46] F. A. Valentine. Some properties of C-convex sets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2(??):481–488, ???? 1950. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Green:1950:QS

[47] J. W. Green and W. Gustin. Quasiconvex sets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2(??):489–507, ???? 1950. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Green:1950:VSC

[48] J. W. Green and W. Gustin. On the vector sum of continua. Canadian Journal of Mathematics = Journal canadien de mathématiques, 2(??):508–512, ???? 1950. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dirac:1951:HFF

[49] P. A. M. Dirac. The Hamiltonian form of field dynamics. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3(??):1–23, ???? 1951.

CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Croisot:1951:ALD

[50] R. Croisot. Axiomatique des lattices distributives. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3(??):24–27, ???? 1951. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sholander:1951:PDL

[51] Marlow Sholander. Postulates for distributive lattices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3(??):28–30, ???? 1951. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Byrne:1951:SFB

[52] Lee Byrne. Short formulations of Boolean algebra, using ring operations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3 (??):31–33, ???? 1951. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lambek:1951:ISG

[53] J. Lambek. The immersibility of a semigroup into a group. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3(??):34–43, ???? 1951. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hua:1951:ESA

[54] Loo-Keng Hua. On exponential sums over an algebraic number field. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3(??): 44–51, ???? 1951. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Porcelli:1951:PFS

[55] Pasquale Porcelli and Gordon Pall. A property of Farey sequences, with applications to qth power residues. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3(??): 52–53, ???? 1951. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Macbeath:1951:CTA

[56] A. M. Macbeath. A compactness theorem for affine equivalence-classes of convex regions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3(??):54–61, ???? 1951. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Burchnall:1951:BP

[57] J. L. Burchnall. The Bessel polynomials. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3(??):62–68, ???? 1951. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Truesdell:1951:VA

[58] C. Truesdell. Vorticity averages. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3(??): 69–86, ???? 1951. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ellis:1951:ABAa

[59] David Ellis. Autometrized Boolean algebras. I. Fundamental distance-theoretic properties of B. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3(??):87–93, ???? 1951. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bruck:1951:FNN

[60] R. H. Bruck. Finite nets, I. Numerical invariants. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3(??):94–107, ???? 1951. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Heins:1951:E

[61] Albert E. Heins. Errata. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3(??):107, ???? 1951. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kodaira:1951:GFM

[62] Kunihiko Kodaira. Green’s forms and meromorphic functions on compact analytic varieties. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3(??):108–128, ???? 1951. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dieudonne:1951:TLN

[63] Jean Dieudonné. Sur le théorème de Lebesgue–Nikodym. V. (French) [On the Lebesgue–Nikodym theorem. V]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3(??): 129–139, ???? 1951. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lukacs:1951:CFT

[64] Eugene Lukacs and Otto Szász. Certain Fourier transforms of distributions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3 (??):140–144, ???? 1951. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ellis:1951:ABAb

[65] David Ellis. Autometrized Boolean algebras. II. The group of motions of B. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3(??):145–147, ???? 1951. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

DuVal:1951:RSG

[66] Patrick Du Val. On regular surfaces of genus three. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3(??):148–154, ???? 1951. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hutcherson:1951:CIP

[67] W. R. Hutcherson. A cyclic involution of period eleven. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3(??):155–158, ???? 1951. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Derry:1951:DTC

[68] Douglas Derry. The duality theorem for curves of order n in n-space. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3(??): 159–163, ???? 1951. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Stanton:1951:MG

[69] R. G. Stanton. The Mathieu groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3 (??):164–174, ???? 1951. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bing:1951:MTS

[70] R. H. Bing. Metrization of topological spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3(??):175–186, ???? 1951. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kelly:1951:DS

[71] L. M. Kelly. Distance sets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3(??): 187–194, ???? 1951. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Infeld:1951:RGE

[72] L. Infeld and A. E. Scheidegger. Radiation and gravitational equations of motion. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3(??):195–207, ???? 1951. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Nakayama:1951:NNG

[73] Tadasi Nakayama. Non-normal Galois theory for non-commutative and non-semisimple rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3(??):208–218, ???? 1951. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dowker:1951:CPS

[74] C. H. Dowker. On countably paracompact spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3(??):219–224, ????

1951. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Steinberg:1951:R

[75] Robert Steinberg. The representations of GL(3, q), GL(4, q), PGL(3, q), and PGL(4, q). Canadian Journal of Mathematics = Journal canadien de mathématiques, 3(??):225–235, ???? 1951. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lorentz:1951:DTM

[76] G. G. Lorentz. Direct theorems on methods of summability. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3(??):236–256, ???? 1951. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Heilbronn:1951:EAC

[77] H. Heilbronn. On Euclid’s algorithm in cyclic fields. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3(??):257–268, ???? 1951. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

ApSimon:1951:TVR

[78] Hugh ApSimon. Two vertex-regular polyhedra. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3(??):269–271, ???? 1951. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Klee:1951:CIP

[79] V. L. Klee, Jr. On certain intersection properties of convex sets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3(??): 272–275, ???? 1951. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Jenkins:1951:TTF

[80] James A. Jenkins. On the topological theory of functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3(??):276–289, ???? 1951. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kaplansky:1951:TDR

[81] Irving Kaplansky. A theorem on division rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3 (??):290–292, ???? 1951. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Halperin:1951:DSO

[82] Israel Halperin and Hidegoro Nakano. Discrete semi-ordered linear spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3 (??):293–298, ???? 1951. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Berman:1951:GPC

[83] Gerald Berman. A generalization of the Pappus configuration. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3(??):299–303, ???? 1951. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Willcocks:1951:NSP

[84] T. H. Willcocks. A note on some perfect squared squares. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3(??):304–308, ???? 1951. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chung:1951:MRS

[85] J. H. Chung. Modular representations of the symmetric group. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3(??):309–327, ???? 1951. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chowla:1951:RCS

[86] S. Chowla, I. N. Herstein, and W. K. Moore. On recursions connected with symmetric groups. I. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3(??):328–334, ???? 1951. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Coddington:1951:NSS

[87] E. A. Coddington and N. Levinson. On the nature of the spectrum of singular second order linear differential equations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3 (??):335–338, ???? 1951. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Reeb:1951:SPC

[88] Georges Reeb. Sur les solutions périodiques de certains systèmes différentiels perturbés. (French) [On the periodic solutions of certain perturbed differential equations]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3(??):339–362, ???? 1951. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fraser:1951:NSL

[89] D. A. S. Fraser. Normal samples with linear constraints and given variances. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3 (??):363–366, ???? 1951. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Brauer:1951:TAT

[90] Alfred Brauer and R. L. Reynolds. On a theorem of Aubry–Thue. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3(??): 367–374, ???? 1951. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Erdos:1951:CSC

[91] Paul Erdős and Harold N. Shapiro. On the changes of sign of a certain error function. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3(??):375–385, ???? 1951. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Macphail:1951:STA

[92] M. S. Macphail. Some theorems on absolute summability. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3(??):386–390, ???? 1951. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Coxeter:1951:EF

[93] H. S. M. Coxeter. Extreme forms. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3(??): 391–441, ???? 1951. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hlavaty:1951:SSL

[94] Václav Hlavatý. Spinor space and line geometry. I. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3(??):442–459, ???? 1951. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bochner:1951:NVD

[95] S. Bochner. A new viewpoint in differential geometry. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3(??):460–470, ???? 1951. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ellis:1951:DPA

[96] H. W. Ellis. Darboux properties and applications to non-absolutely convergent integrals. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3(??):471–485, ???? 1951. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ankeny:1951:CNC

[97] N. C. Ankeny and S. Chowla. The class number of the cyclotomic field. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3(??): 486–494, ???? 1951. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hall:1951:CIM

[98] Marshall Hall and H. J. Ryser. Cyclic incidence matrices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 3(??):495–502, ???? 1951. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Schwartz:1951:ASH

[99] Laurent Schwartz. Analyse et synthèse harmoniques dans les espaces de distributions. (French) [Harmonic analysis and synthesis in distribution space].

Canadian Journal of Mathematics = Journal canadien de mathématiques, 3 (??):503–512, ???? 1951. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Touchard:1952:PCF

[100] Jacques Touchard. Sur un problème de configurations et sur les fractions continues. (French) [On a configuration problem and on continued fractions]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 4 (??):2–25, ???? 1952. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Minakshisundaram:1952:ZFU

[101] S. Minakshisundaram. Zeta functions on the unitary sphere. Canadian Journal of Mathematics = Journal canadien de mathématiques, 4(??):26–30, ???? 1952. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Goldhaber:1952:HMC

[102] J. K. Goldhaber. The homomorphic mapping of certain matric algebras onto rings of diagonal matrices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 4(??): 31–42, ???? 1952. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Murdoch:1952:CNI

[103] D. C. Murdoch. Contributions to noncommutative ideal theory. Canadian Journal of Mathematics = Journal canadien de mathématiques, 4(??): 43–57, ???? 1952. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Davenport:1952:NND

[104] H. Davenport and P. Erdős. Note on normal decimals. Canadian Journal of Mathematics = Journal canadien de mathématiques, 4(??):58–63, ???? 1952. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mann:1952:PSG

[105] Henry B. Mann. On products of sets of group elements. Canadian Journal of Mathematics = Journal canadien de mathématiques, 4(??):64–66, ???? 1952. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Simons:1952:FCM

[106] William H. Simons. The Fourier coefficients of the modular function λ(τ). Canadian Journal of Mathematics = Journal canadien de mathématiques, 4 (??):67–80, ???? 1952. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gans:1952:AEG

[107] David Gans. Axioms for elliptic geometry. Canadian Journal of Mathematics = Journal canadien de mathématiques, 4(??):81–92, ???? 1952. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Feld:1952:GLE

[108] J. M. Feld. On the geometry of lineal elements on a sphere, Euclidean kinematics, and elliptic geometry. Canadian Journal of Mathematics = Journal canadien de mathématiques, 4(??): 93–110, ???? 1952. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rothberger:1952:PPH

[109] Fritz Rothberger. On the property C and a problem of Hausdorff. Canadian Journal of Mathematics = Journal canadien de mathématiques, 4(??):111–116, ???? 1952. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rothberger:1952:RED

[110] Fritz Rothberger. A remark on the existence of a denumerable base for a family of functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 4(??):117–119, ???? 1952. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Jones:1952:EMT

[111] Burton W. Jones. An extension of Meyer’s theorem on indefinite ternary quadratic forms. Canadian Journal of Mathematics = Journal canadien de mathématiques, 4(??):120–128, ???? 1952. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See correction [174].

Albert:1952:SAR

[112] A. A. Albert. On simple alternative rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 4 (??):129–135, ???? 1952. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Beatty:1952:SPR

[113] S. Beatty and N. D. Lane. A symmetric proof of the Riemann–Roch theorem, and a new form of the unit theorem. Canadian Journal of Mathematics = Journal canadien de mathématiques, 4(??):136–148, ???? 1952. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fox:1952:IT

[114] Charles Fox. Iterated transforms. Canadian Journal of Mathematics = Journal canadien de mathématiques, 4(??): 149–161, ???? 1952. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ayoub:1952:TNC

[115] Christine Williams Ayoub. A theory of normal chains. Canadian Journal of Mathematics = Journal canadien de mathématiques, 4(??):162–188, ???? 1952. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Moor:1952:QRG

[116] Arthur Moór. Quelques remarques sur la généralisation du scalaire de courbure et du scalaire principal. (French) [Some remarks on the generalization of the scalar curvature and the principal scalar]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 4 (??):189–197, ???? 1952. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Graves:1952:CCO

[117] Ross E. Graves. A closure criterion for orthogonal functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 4(??):198–203, ???? 1952. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

DuVal:1952:SWC

[118] Patrick Du Val. On surfaces whose canonical system is hyperelliptic. Canadian Journal of Mathematics = Journal canadien de mathématiques, 4(??): 204–221, ???? 1952. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mann:1952:STD

[119] Henry B. Mann. Some theorems on difference sets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 4(??):222–226, ???? 1952. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Thrall:1952:GCB

[120] R. M. Thrall. On a Galois connection between algebras of linear transformations and lattices of subspaces of a vector space. Canadian Journal of Mathematics = Journal canadien de mathématiques, 4(??):227–239, ???? 1952. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Frucht:1952:ORG

[121] Robert Frucht. A one-regular graph of degree three. Canadian Journal of Mathematics = Journal canadien de mathématiques, 4(??):240–247, ???? 1952. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Duff:1952:EFT

[122] G. F. D. Duff. F -equation Fourier transforms. Canadian Journal of Mathematics = Journal canadien de mathématiques, 4(??):248–256, ???? 1952. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Karush:1952:IPC

[123] William Karush. Isoperimetric problems in the calculus of variations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 4(??): 257–280, ???? 1952. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Karush:1952:IEA

[124] William Karush. The index of an extremal arc. Canadian Journal of Mathematics = Journal canadien de mathématiques, 4(??):281–294, ???? 1952. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hoffman:1952:CAP

[125] A. J. Hoffman. Cyclic affine planes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 4 (??):295–301, ???? 1952. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Berman:1952:FPG

[126] Gerald Berman. Finite projective geometries. Canadian Journal of Mathematics = Journal canadien de mathématiques, 4(??):302–313, ???? 1952. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Tutte:1952:FG

[127] W. T. Tutte. The factors of graphs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 4 (??):314–328, ???? 1952. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

White:1952:SCG

[128] Paul A. White. Some characterizations of generalized manifolds with boundaries. Canadian Journal of Mathematics = Journal canadien de mathématiques, 4(??):329–342, ???? 1952. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Faircloth:1952:NSS

[129] Olin B. Faircloth. On the number of solutions of some general types of equations in a finite field. Canadian Journal of Mathematics = Journal canadien de mathématiques, 4(??):343–351, ???? 1952. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ramanathan:1952:AQF

[130] K. G. Ramanathan. Abelian quadratic forms. Canadian Journal of Mathematics = Journal canadien de mathématiques, 4(??):352–368, ???? 1952. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Haimo:1952:GCC

[131] Franklin Haimo. Groups with a certain condition on conjugates. Canadian Journal of Mathematics = Journal canadien de mathématiques, 4(??): 369–372, ???? 1952. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Robinson:1952:CJH

[132] G. de B. Robinson. On a conjecture by J. H. Chung. Canadian Journal of Mathematics = Journal canadien de mathématiques, 4(??):373–380, ???? 1952. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Osima:1952:IRS

[133] Masaru Osima. On the irreducible representations of the symmetric group. Canadian Journal of Mathematics = Journal canadien de mathématiques, 4 (??):381–384, ???? 1952. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Stevenson:1952:PFD

[134] A. F. Stevenson and W. A. Bassali. On the possible forms of differential equation which can be factorized by the Schrödinger–Infeld method. Canadian Journal of Mathematics = Journal canadien de mathématiques, 4(??): 385–395, ???? 1952. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gelbaum:1952:MS

[135] B. R. Gelbaum and G. K. Kalisch. Measure in semigroups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 4(??):396–406, ???? 1952. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Robertson:1952:CPF

[136] M. S. Robertson. A coefficient problem for functions regular in an annulus. Canadian Journal of Mathematics = Journal canadien de mathématiques, 4(??):407–423, ???? 1952. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Glover:1952:AFI

[137] I. E. Glover. Analytic functions with an irregular linearly measurable set of singular points. Canadian Journal of Mathematics = Journal canadien de mathématiques, 4(??):424–435, ???? 1952. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rooney:1952:NRI

[138] P. G. Rooney. A new representation and inversion theory for the Laplace transformation. Canadian Journal of Mathematics = Journal canadien de mathématiques, 4(??):436–444, ???? 1952. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Petersen:1952:NDS

[139] G. M. Petersen. A note on divergent series. Canadian Journal of Mathematics = Journal canadien de mathématiques, 4(??):445–454, ???? 1952. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lorentz:1952:MSF

[140] G. G. Lorentz. Multiply subadditive functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 4(??):455–462, ???? 1952. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Halperin:1952:SFA

[141] Israel Halperin. The supremum of a family of additive functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 4(??):463–479, ???? 1952. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dirac:1952:MCT

[142] G. A. Dirac. Map-colour theorems.

Canadian Journal of Mathematics = Journal canadien de mathématiques, 4 (??):480–490, ???? 1952. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Zemmer:1952:SFD

[143] Joseph L. Zemmer, Jr. On the subalgebras of finite division algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 4(??): 491–503, ???? 1952. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Duncan:1952:DLA

[144] D. G. Duncan. On D. E. Littlewood’s algebra of S-functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 4(??):504–512, ???? 1952. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Matte:1953:NSO

[145] Alphonse Matte. Sur de nouvelles solutions oscillatoires de équations de la gravitation. (French) [On new oscillatory solutions of the gravitational equations]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??):1–16, ???? 1953. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Infeld:1953:CCE

[146] L. Infeld. The coordinate conditions and the equations of motion. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??):17–25, ???? 1953. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Scheidegger:1953:MQ

[147] A. E. Scheidegger. Multiple quantization. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??):26–36, ???? 1953. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Martin:1953:PPS

[148] M. H. Martin. The propagation of a plane shock into a quiet atmosphere. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5 (??):37–39, ???? 1953. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Roeser:1953:CPT

[149] Ernst Roeser. Corresponding polyhedra in the three spaces of constant curvature. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??):40–45, ???? 1953. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Synge:1953:SBF

[150] J. L. Synge. A simple bounding formula for integrals. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??):46–52, ???? 1953. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hawley:1953:CHC

[151] N. S. Hawley. Constant holomorphic curvature. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??):53–56, ???? 1953. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Duff:1953:HTN

[152] G. F. D. Duff. Harmonic p-tensors on normal hyperbolic Riemannian spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5 (??):57–80, ???? 1953. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gustin:1953:PAI

[153] William Gustin. Partitioning an arithmetic interval. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??):81–85, ???? 1953. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Edrei:1953:PCS

[154] Albert Edrei. Proof of a conjecture of Schoenberg on the generating function of a totally positive sequence. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??): 86–94, ???? 1953. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Pierce:1953:BAR

[155] R. S. Pierce. The Boolean algebra of regular open sets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??):95–100, ???? 1953. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ewing:1953:CSF

[156] G. M. Ewing and W. R. Utz. Continuous solutions of the functional equation f^{n}(x) = f(x). Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??):101–103, ???? 1953. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rainville:1953:GFB

[157] E. D. Rainville. Generating functions for Bessel and related polynomials. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??): 104–106, ???? 1953. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Butzer:1953:TDB

[158] P. L. Butzer. On two-dimensional Bernstein polynomials. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??):107–113, ???? 1953. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sumner:1953:CTA

[159] D. B. Sumner. A convolution transform admitting an inversion formula of integro-differential type. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??): 114–117, ???? 1953. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hirschman:1953:SPD

[160] I. I. Hirschman, Jr. Systems of partial differential equations which generalize the heat equation. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??):118–128, ???? 1953. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Grothendieck:1953:ALF

[161] A. Grothendieck. Sur les applications linéaires faiblement compactes d’espaces du type C(K). (French) [On the applications of linear weakly-compact spaces of the type C(K)]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??):129–173, ???? 1953. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bourgin:1953:SMF

[162] D. G. Bourgin. Some multiplicative functionals. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??):174–178, ???? 1953. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Boas:1953:IFN

[163] R. P. Boas, Jr. Integral functions with negative zeros. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??):179–184, ???? 1953. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Slepian:1953:NST

[164] David Slepian. On the number of symmetry types of Boolean functions of n variables. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??):185–193, ???? 1953. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rudin:1953:RCG

[165] Walter Rudin. A remark concerning Graves’ closure criterion. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??): 194–195, ???? 1953. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Duff:1953:BVP

[166] G. F. D. Duff. Boundary value problems associated with the tensor Laplace equation. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??):196–210, ???? 1953. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Arens:1953:ECP

[167] Richard Arens. Extension of coverings, of pseudometrics, and of linearspace-valued mappings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??):211–215, ???? 1953. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

DuVal:1953:RSG

[168] Patrick Du Val. Regular surfaces of genus two. I. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??):216–237, ???? 1953. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Herstein:1953:TR

[169] I. N. Herstein. A theorem on rings.

Canadian Journal of Mathematics = Journal canadien de mathématiques, 5 (??):238–241, ???? 1953. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Nakayama:1953:CCD

[170] Tadasi Nakayama. On the commutativity of certain division rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??): 242–244, ???? 1953. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Moser:1953:NAS

[171] Leo Moser. On non-averaging sets of integers. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??):245–252, ???? 1953. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Churchhouse:1953:CI

[172] R. F. Churchhouse. A criterion for irrationality. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??):253–260, ????

1953. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cohn:1953:SL

[173] Harvey Cohn. Stable lattices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??): 261–270, ???? 1953. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Jones:1953:CEM

[174] B. W. Jones. Correction to “An extension of Meyer’s theorem on indefinite ternary quadratic forms”. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??): 271–272, ???? 1953. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See [111].

Halperin:1953:FS

[175] Israel Halperin. Function spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??): 273–288, ???? 1953. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hill:1953:SMD

[176] J. D. Hill. Summability methods defined by Riemann sums. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??):289–296, ???? 1953. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lamson:1953:CTE

[177] K. W. Lamson. On the curvature tensor of Einstein’s generalized theory of gravitation. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??):297–300, ????

1953. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Brafman:1953:RBU

[178] Fred Brafman. A relation between ultraspherical and Jacobi polynomial sets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5 (??):301–305, ???? 1953. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Carlitz:1953:TG

[179] Leonard Carlitz. A theorem of Glaisher.

Canadian Journal of Mathematics = Journal canadien de mathématiques, 5 (??):306–316, ???? 1953. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Carlitz:1953:WQP

[180] Leonard Carlitz. Weighted quadratic partitions over a finite field. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??): 317–323, ???? 1953. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Elliott:1953:ASD

[181] J. G. Elliott. Autometrization and the symmetric difference. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??):324–331, ???? 1953. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Goldhaber:1953:SMT

[182] J. K. Goldhaber and G. Whaples.

On some matrix theorems of Frobenius and McCoy. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??):332–335, ????

1953. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Osima:1953:SRC

[183] Masaru Osima. Some remarks on the characters of the symmetric group. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5 (??):336–343, ???? 1953. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Maranda:1953:AIR

[184] Jean-Marie Maranda. On B-adic integral representations of finite groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5 (??):344–355, ???? 1953. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Nagao:1953:NMR

[185] Hirosi Nagao. Note on the modular representations of symmetric groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??): 356–363, ???? 1953. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Shephard:1953:UGG

[186] G. C. Shephard. Unitary groups generated by reflections. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??):364–383, ???? 1953. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Coxeter:1953:EDF

[187] H. S. M. Coxeter and J. A. Todd. An extreme duodenary form. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??):384–392, ????

1953. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Tyler:1953:NIF

[188] G. W. Tyler. Numerical integration of functions of several variables. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??): 393–412, ???? 1953. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Shrikhande:1953:NEC

[189] S. S. Shrikhande. The non-existence of certain affine resolvable balanced incomplete block designs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??):413–420, ???? 1953. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ostrom:1953:CDS

[190] T. G. Ostrom. Concerning difference sets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??):421–424, ???? 1953. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lehmer:1953:RDS

[191] Emma Lehmer. On residue difference sets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??):425–432, ???? 1953. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mordell:1953:NLS

[192] L. J. Mordell. Note on the linear symmetric congruence in n variables. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??):

433–438, ???? 1953. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ayoub:1953:WST

[193] R. G. Ayoub. On the Waring–Siegel theorem. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5 (??):439–450, ???? 1953. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Watson:1953:IRP

[194] G. L. Watson. On integers n relatively prime to [αn]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??):451–455, ???? 1953. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Estermann:1953:NPL

[195] Theodor Estermann. On the number of primitive lattice points in a parallelogram. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5 (??):456–459, ???? 1953. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sholander:1953:PBA

[196] Marlow Sholander. Postulates for Boolean algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??):460–464, ???? 1953. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Copeland:1953:CIB

[197] A. H. Copeland, Sr. and Frank Harary. A characterization of implicative Boolean rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??):465–469, ????

1953. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Goldberg:1953:ELA

[198] S. I. Goldberg. Extensions of Lie algebras and the third cohomology group. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5 (??):470–476, ???? 1953. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Higman:1953:FSF

[199] D. G. Higman. Focal series in finite groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??):477–497, ???? 1953. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Haimo:1953:CG

[200] Franklin Haimo. The F C-chain of a group. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??):498–511, ???? 1953. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lane:1953:DPC

[201] N. D. Lane and Peter Scherk. Differentiable points in the conformal plane. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5 (??):512–518, ???? 1953. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Pan:1953:VCC

[202] T. K. Pan. Variation of congruences of curves of an orthogonal ennuple in a Riemannian space. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??):519–523, ????

1953. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Duff:1953:TEE

[203] G. F. D. Duff. A tensor equation of elliptic type. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??):524–535, ???? 1953. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Walmsley:1953:NTS

[204] Charles Walmsley. Null trigonometric series in differential equations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??): 536–543, ???? 1953. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See correction [254].

Agarwal:1953:GTB

[205] R. P. Agarwal. General transformations of bilateral cognate trigonometrical series of ordinary hypergeometric type. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??):544–553, ???? 1953. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dyson:1953:FTD

[206] F. J. Dyson. Fourier transforms of distribution functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??):554–558, ???? 1953. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Butzer:1953:LCB

[207] P. L. Butzer. Linear combinations of Bernstein polynomials. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??):559–567, ????

1953. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lorentz:1953:RLF

[208] G. G. Lorentz and D. G. Wertheim.

Representation of linear functionals on Köthe spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??):568–575, ???? 1953. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ellis:1953:FSD

[209] H. W. Ellis and Israel Halperin. Function spaces determined by a levelling length function. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??):576–592, ???? 1953. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Brouwer:1954:PS

[210] L. E. J. Brouwer. Points and spaces.

Canadian Journal of Mathematics = Journal canadien de mathématiques, 6 (??):1–17, ???? 1954. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

vanderPol:1954:NGF

[211] B. van der Pol. Note on the gamma function. Canadian Journal of Mathematics = Journal canadien de mathématiques, 6(??):18–22, ???? 1954. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Carlitz:1954:FFC

[212] Leonard Carlitz. The first factor of the class number of a cyclic field. Canadian Journal of Mathematics = Journal canadien de mathématiques, 6(??):

23–26, ???? 1954. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lorentz:1954:CCT

[213] G. G. Lorentz and A. Robinson. Coreconsistency and total inclusion for methods of summability. Canadian Journal of Mathematics = Journal canadien de mathématiques, 6(??):27–34, ???? 1954. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hall:1954:ETB

[214] Marshall Hall, Jr. and W. S. Connor. An embedding theorem for balanced incomplete block designs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 6(??):35–41, ???? 1954. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fraser:1954:COS

[215] D. A. S. Fraser. Completeness of order statistics. Canadian Journal of Mathematics = Journal canadien de mathématiques, 6(??):42–45, ???? 1954. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fraser:1954:NPT

[216] D. A. S. Fraser. Non-parametric theory. Canadian Journal of Mathematics = Journal canadien de mathématiques, 6(??):46–68, ???? 1954. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mitchell:1954:UT

[217] B. E. Mitchell. Unitary transformations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 6(??):69–72, ???? 1954. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Duff:1954:LPD

[218] G. F. D. Duff. On linear partial differential equations of the second order having geodesic solutions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 6(??):73–79, ???? 1954. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Tutte:1954:CTC

[219] W. T. Tutte. A contribution to the theory of chromatic polynomials. Canadian Journal of Mathematics = Journal canadien de mathématiques, 6(??): 80–91, ???? 1954. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hutcherson:1954:MCA

[220] W. R. Hutcherson and S. T. Gormsen.

Maps of certain algebraic curves invariant under cyclic involutions of periods three, five, and seven. Canadian Journal of Mathematics = Journal canadien de mathématiques, 6(??):92–98, ???? 1954. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Henderson:1954:PC

[221] G. P. Henderson. Parallel curves. Canadian Journal of Mathematics = Journal canadien de mathématiques, 6(??): 99–107, ???? 1954. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Behrend:1954:SGT

[222] F. A. Behrend. The Steinitz–Gross theorem on sums of vectors. Canadian Journal of Mathematics = Journal canadien de mathématiques, 6(??):108–124, ???? 1954. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Frame:1954:TON

[223] J. S. Frame and G. de B. Robinson. On a theorem of Osima and Nagao. Canadian Journal of Mathematics = Journal canadien de mathématiques, 6(??): 125–127, ???? 1954. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Shephard:1954:CWP

[224] G. C. Shephard. A construction for Wythoffian polytopes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 6(??):128–134, ???? 1954. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Meyer:1954:SSH

[225] Burnett Meyer. On the symmetries of spherical harmonics. Canadian Journal of Mathematics = Journal canadien de mathématiques, 6(??):135–157, ???? 1954. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Grothendieck:1954:CSE

[226] A. Grothendieck. Sur certains sousespaces vectoriels de L^{p}. Canadian Journal of Mathematics = Journal canadien de mathématiques, 6(??):158–160, ???? 1954. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

McShane:1954:TC

[227] E. J. McShane. A theory of convergence. Canadian Journal of Mathematics = Journal canadien de mathématiques, 6 (??):161–168, ???? 1954. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Coddington:1954:SMG

[228] E. A. Coddington. The spectral matrix and Green’s function for singular selfadjoint boundary value problems. Canadian Journal of Mathematics = Journal canadien de mathématiques, 6(??): 169–185, ???? 1954. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lukacs:1954:CFT

[229] Eugene Lukacs and Otto Szász. Certain Fourier transforms of distributions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 6 (??):186–189, ???? 1954. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rooney:1954:IRT

[230] P. G. Rooney. An inversion and representation theory for the Laplace integral of abstractly-valued functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 6(??): 190–209, ???? 1954. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Pan:1954:GFC

[231] T. K. Pan. On a generalization of the first curvature of a curve in a hypersurface of a Riemannian space. Canadian Journal of Mathematics = Journal canadien de mathématiques, 6(??): 210–216, ???? 1954. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See correction [289].

Nordhaus:1954:BG

[232] E. A. Nordhaus and L. Lapidus. Brouwerian geometry. Canadian Journal of Mathematics = Journal canadien de mathématiques, 6(??):217–229, ???? 1954. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Scherk:1954:CAG

[233] Peter Scherk and J. H. B. Kemperman. Complexes in Abelian groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 6(??): 230–237, ???? 1954. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kemperman:1954:SSI

[234] J. H. B. Kemperman and Peter Scherk. On sums of sets of integers. Canadian Journal of Mathematics = Journal canadien de mathématiques, 6(??): 238–252, ???? 1954. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cohn:1954:HIS

[235] P. M. Cohn. On homomorphic images of special Jordan algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 6(??):253–264, ???? 1954. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cohn:1954:SL

[236] Harvey Cohn. Stable lattices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 6(??): 265–273, ???? 1954. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Shephard:1954:FUR

[237] G. C. Shephard and J. A. Todd. Finite unitary reflection groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 6(??): 274–304, ???? 1954. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dieudonne:1954:IEE

[238] Jean Dieudonné. Les isomorphismes exceptionnels entre les groupes classiques finis. Canadian Journal of Mathematics = Journal canadien de mathématiques, 6(??):305–315, ???? 1954. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Frame:1954:HGS

[239] J. S. Frame, G. de B. Robinson, and R. M. Thrall. The hook graphs of the symmetric group. Canadian Journal of Mathematics = Journal canadien de mathématiques, 6(??):316–324, ???? 1954. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Jennings:1954:SGF

[240] S. A. Jennings. Substitution groups of formal power series. Canadian Journal of Mathematics = Journal canadien de mathématiques, 6(??):325–340, ???? 1954. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sprott:1954:NBI

[241] D. A. Sprott. A note on balanced incomplete block designs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 6(??):341–346, ???? 1954. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Tutte:1954:SPF

[242] W. T. Tutte. A short proof of the factor theorem for finite graphs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 6(??): 347–352, ???? 1954. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Briggs:1954:EPT

[243] W. E. Briggs. An elementary proof of a theorem about the representation of primes by quadratic forms. Canadian Journal of Mathematics = Journal canadien de mathématiques, 6(??): 353–363, ???? 1954. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Whiteman:1954:SPR

[244] A. L. Whiteman. The sixteenth power residue character of 2. Canadian Journal of Mathematics = Journal canadien de mathématiques, 6(??):364–373, ???? 1954. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kaplansky:1954:RIB

[245] Irving Kaplansky. Ring isomorphisms of Banach algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 6(??):374–381, ???? 1954. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Agmon:1954:RML

[246] Shmuel Agmon. The relaxation method for linear inequalities. Canadian Journal of Mathematics = Journal canadien de mathématiques, 6(3):382–392, ???? 1954. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Motzkin:1954:RML

[247] T. S. Motzkin and I. J. Schoenberg. The relaxation method for linear inequalities. Canadian Journal of Mathematics = Journal canadien de mathématiques, 6(3):393–404, ???? 1954. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Macphail:1954:SRD

[248] M. S. Macphail. On some recent developments in the theory of series. Canadian Journal of Mathematics = Journal canadien de mathématiques, 6(??): 405–409, ???? 1954. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Agarwal:1954:BP

[249] R. P. Agarwal. On Bessel polynomials. Canadian Journal of Mathematics = Journal canadien de mathématiques, 6(??):410–415, ???? 1954. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sternberg:1954:TPB

[250] H. M. Sternberg and R. L. Sternberg. A two-point boundary problem for ordinary self-adjoint differential equations of fourth order. Canadian Journal of Mathematics = Journal canadien de mathématiques, 6(??):416–419, ???? 1954. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Putnam:1954:CSS

[251] C. R. Putnam. On the continuous spectra of singular boundary value problems. Canadian Journal of Mathematics = Journal canadien de mathématiques, 6(??):420–426, ???? 1954. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Duff:1954:TBV

[252] G. F. D. Duff. A tensor boundary value problem of mixed type. Canadian Journal of Mathematics = Journal canadien de mathématiques, 6(??):427–440, ???? 1954. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Garabedian:1954:DIV

[253] P. R. Garabedian and M. Schiffer. On a double integral variational problem. Canadian Journal of Mathematics = Journal canadien de mathématiques, 6 (??):441–446, ???? 1954. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Walmsley:1954:CNT

[254] Charles Walmsley. Correction to “Null trigonometric series in differential equations”. Canadian Journal of Mathematics = Journal canadien de mathématiques, 6(??):447–448, ???? 1954. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See [204].

Lehmer:1954:CNO

[255] Emma Lehmer. On the cyclotomic numbers of order sixteen. Canadian Journal of Mathematics = Journal canadien de mathématiques, 6(??):449–454, ???? 1954. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ward:1954:MPD

[256] Morgan Ward. The maximal prime divisors of linear recurrences. Canadian Journal of Mathematics = Journal canadien de mathématiques, 6(??): 455–462, ???? 1954. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chowla:1954:DBQ

[257] S. Chowla and W. E. Briggs. On discriminants of binary quadratic forms with a single class in each genus. Canadian Journal of Mathematics = Journal canadien de mathématiques, 6(??): 463–470, ???? 1954. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Butts:1954:IC

[258] Hubert Butts, Marshall Hall, Jr., and H. B. Mann. On integral closure. Canadian Journal of Mathematics = Journal canadien de mathématiques, 6(??): 471–473, ???? 1954. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mann:1954:EPC

[259] H. B. Mann. On an exceptional phenomenon in certain quadratic extensions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 6 (??):474–476, ???? 1954. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hanson:1954:SRB

[260] H. A. Hanson. Some relations between various types of normality of numbers. Canadian Journal of Mathematics = Journal canadien de mathématiques, 6 (??):477–485, ???? 1954. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Robinson:1954:MRS

[261] G. de B. Robinson. On the modular representations of the symmetric group Part IV. Canadian Journal of Mathematics = Journal canadien de mathématiques, 6(??):486–497, ???? 1954. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Burrow:1954:GYD

[262] M. D. Burrow. A generalization of the Young diagram. Canadian Journal of Mathematics = Journal canadien de mathématiques, 6(??):498–508, ???? 1954. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Duncan:1954:NAF

[263] D. G. Duncan. Note on the algebra of S-functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 6(??):509–510, ???? 1954. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Osima:1954:SRC

[264] Masaru Osima. Some remarks on the characters of the symmetric group. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 6 (??):511–521, ???? 1954. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hamilton:1954:SPC

[265] O. H. Hamilton. A short proof of the Cartwright–Littlewood fixed point theorem. Canadian Journal of Mathematics = Journal canadien de mathématiques, 6(??):522–524, ???? 1954. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Botts:1954:LEP

[266] Truman Botts. On lattice embeddings for partially ordered sets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 6(??): 525–528, ???? 1954. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Barrett:1954:DEN

[267] J. H. Barrett. Differential equations of non-integer order. Canadian Journal of Mathematics = Journal canadien de mathématiques, 6(??):529–541, ???? 1954. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Protter:1954:CPH

[268] M. H. Protter. The Cauchy problem for a hyperbolic second order equation with data on the parabolic line. Canadian Journal of Mathematics = Journal canadien de mathématiques, 6(??): 542–553, ???? 1954. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Conte:1954:ETP

[269] S. D. Conte and W. C. Sangren. An expansion theorem for a pair of singular first order equations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 6(??):554–560, ???? 1954. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Atkinson:1954:LPN

[270] F. V. Atkinson. On linear perturbation of non-linear differential equations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 6 (??):561–571, ???? 1954. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Butzer:1954:RDI

[271] P. L. Butzer and W. Kozakiewicz. On the Riemann derivatives for integrable functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 6(??):572–581, ???? 1954. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Shapiro:1954:LCS

[272] V. L. Shapiro. Logarithmic capacity of sets and double trigonometric series. Canadian Journal of Mathematics = Journal canadien de mathématiques, 6(??):582–592, ???? 1954. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Greenwood:1955:CRC

[273] R. E. Greenwood and A. M. Gleason. Combinatorial relations and chromatic graphs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 7(??):1–7, ???? 1955. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Frucht:1955:RFG

[274] Robert Frucht. Remarks on finite groups defined by generating relations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 7 (??):8–17, ???? 1955. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Coxeter:1955:LGG

[275] H. S. M. Coxeter. On Laves’ graph of girth ten. Canadian Journal of Mathematics = Journal canadien de mathématiques, 7(??):18–23, ???? 1955. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fryer:1955:CPG

[276] K. D. Fryer. A class of permutation groups of prime degree. Canadian Journal of Mathematics = Journal canadien de mathématiques, 7(??):24–34, ???? 1955. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Beaumont:1955:STP

[277] R. A. Beaumont and R. P. Peterson. Settransitive permutation groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 7(??): 35–42, ???? 1955. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Eilenberg:1955:HTA

[278] Samuel Eilenberg and Saunders MacLane.

On the homology theory of abelian groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 7(??):43–53, ???? 1955. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fuller:1955:CSM

[279] L. E. Fuller. A canonical set for matrices over a principal ideal ring modulo m. Canadian Journal of Mathematics = Journal canadien de mathématiques, 7(??):54–59, ???? 1955. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Goheen:1955:WT

[280] Harry Goheen. The Wedderburn theorem. Canadian Journal of Mathematics = Journal canadien de mathématiques, 7(??):60–62, ???? 1955. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Erdos:1955:EDF

[281] Paul Erdős and H. N. Shapiro. The existence of a distribution function for an error term related to the Euler function. Canadian Journal of Mathematics = Journal canadien de mathématiques, 7(??):63–75, ???? 1955. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Jenkins:1955:EBS

[282] J. A. Jenkins. On explicit bounds in Schottky’s theorem. Canadian Journal of Mathematics = Journal canadien de mathématiques, 7(??):76–82, ???? 1955. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Goodman:1955:ROU

[283] A. W. Goodman and E. Reich. On regions omitted by univalent functions. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 7(??):83–88, ???? 1955. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sumner:1955:CTR

[284] D. B. Sumner. Convolution transforms related to nonharmonic Fourier series. Canadian Journal of Mathematics = Journal canadien de mathématiques, 7 (??):89–100, ???? 1955. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rooney:1955:IFL

[285] P. G. Rooney. On an inversion formula for the Laplace transformation. Canadian Journal of Mathematics = Journal canadien de mathématiques, 7(??): 101–115, ???? 1955. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hull:1955:ABI

[286] T. E. Hull and C. Froese. Asymptotic behaviour of the inverse of a Laplace transform. Canadian Journal of Mathematics = Journal canadien de mathématiques, 7(??):116–125, ???? 1955. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Duff:1955:PTC

[287] G. F. D. Duff. On the potential theory of coclosed harmonic forms. Canadian Journal of Mathematics = Journal canadien de mathématiques, 7(??): 126–137, ???? 1955. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ayoub:1955:SLA

[288] R. G. Ayoub. On Selberg’s lemma for algebraic fields. Canadian Journal of Mathematics = Journal canadien de mathématiques, 7(??):138–143, ???? 1955. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Pan:1955:CGF

[289] T. K. Pan. Correction to “On a generalization of the first curvature of a curve in a hypersurface of a Riemannian space”. Canadian Journal of Mathematics = Journal canadien de mathématiques, 7 (??):144, ???? 1955. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See [231].

Kneser:1955:TRE

[290] Martin Kneser. Two remarks on extreme forms. Canadian Journal of Mathematics = Journal canadien de mathématiques, 7(??):145–149, ???? 1955. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Barnes:1955:NEF

[291] E. S. Barnes. Note on extreme forms. Canadian Journal of Mathematics = Journal canadien de mathématiques, 7 (??):150–154, ???? 1955. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lambek:1955:IRP

[292] Joachim Lambek and Leo Moser. On integers n relatively prime to f(n). Canadian Journal of Mathematics = Journal canadien de mathématiques, 7(??): 155–158, ???? 1955. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Moser:1955:SSG

[293] Leo Moser and Max Wyman. On solutions of xspd = 1 in symmetric groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 7 (??):159–168, ???? 1955. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Jennings:1955:GRC

[294] S. A. Jennings. The group ring of a class of infinite nilpotent groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 7(??):

169–187, ???? 1955. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Brown:1955:GMA

[295] W. P. Brown. Generalized matrix algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 7(??):188–190, ???? 1955. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wiegmann:1955:STM

[296] N. A. Wiegmann. Some theorems on matrices with real quaternion elements. Canadian Journal of Mathematics = Journal canadien de mathématiques, 7 (??):191–201, ???? 1955. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Herstein:1955:TCT

[297] I. N. Herstein. A theorem concerning three fields. Canadian Journal of Mathematics = Journal canadien de mathématiques, 7(??):202–203, ???? 1955. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Morelock:1955:AST

[298] J. C. Morelock and N. C. Perry. On algebraic surfaces termwise invariant under cyclic collineations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 7(??):204–207, ???? 1955. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Montgomery:1955:TFP

[299] Deane Montgomery and Hans Samelson. A theorem on fixed points of involutions in Ssp3. Canadian Journal of Mathematics = Journal canadien de mathématiques, 7(??):208–220, ???? 1955. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hayes:1955:FIC

[300] C. A. Hayes, Jr. and C. Y. Pauc. Full individual and class differentiation theorems in their relations to halo and Vitali properties. Canadian Journal of Mathematics = Journal canadien de mathématiques, 7(??):221–274, ???? 1955. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Michael:1955:PFL

[301] Ernest Michael. Point-finite and locally finite coverings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 7(??):275–279, ???? 1955. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Putnam:1955:DSO

[302] C. R. Putnam. On dynamical systems with one degree of freedom. Canadian Journal of Mathematics = Journal canadien de mathématiques, 7(??): 280–283, ???? 1955. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Martin:1955:PPS

[303] M. H. Martin. The propagation of a plane shock into a quiet atmosphere. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 7(??):284–288, ???? 1955. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bartle:1955:WCV

[304] R. G. Bartle, N. Dunford, and J. Schwartz. Weak compactness and vector measures. Canadian Journal of Mathematics = Journal canadien de mathématiques, 7(??):289–305, ???? 1955. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Helfenstein:1955:CML

[305] H. G. Helfenstein. Conformal maps with least distortion. Canadian Journal of Mathematics = Journal canadien de mathématiques, 7(??):306–313, ???? 1955. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rooney:1955:ASS

[306] P. G. Rooney. An application of some spaces of Lorentz. Canadian Journal of Mathematics = Journal canadien de mathématiques, 7(??):314–321, ???? 1955. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kadison:1955:IFI

[307] R. V. Kadison. Isomorphisms of factors of infinite type. Canadian Journal of Mathematics = Journal canadien de mathématiques, 7(??):322–327, ???? 1955. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rosenthall:1955:RDE

[308] E. Rosenthall. Reducible diophantine equations and their parametric representations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 7(??):328–336, ???? 1955. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bambah:1955:IMN

[309] R. P. Bambah and K. Rogers. An inhomogeneous minimum for nonconvex star-regions with hexagonal symmetry.

Canadian Journal of Mathematics = Journal canadien de mathématiques, 7 (??):337–346, ???? 1955. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lehmer:1955:DT

[310] D. H. Lehmer. The distribution of totatives. Canadian Journal of Mathematics = Journal canadien de mathématiques, 7(??):347–357, ???? 1955. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Butson:1955:SLC

[311] A. T. Butson and B. M. Stewart. Systems of linear congruences. Canadian Journal of Mathematics = Journal canadien de mathématiques, 7(??): 358–368, ???? 1955. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sprott:1955:SSP

[312] D. A. Sprott. Some series of partially balanced incomplete block designs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 7 (??):369–381, ???? 1955. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Thrall:1955:CAU

[313] R. M. Thrall. A class of algebras without unity element. Canadian Journal of Mathematics = Journal canadien de mathématiques, 7(??):382–390, ???? 1955. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Robinson:1955:MRS

[314] G. de B. Robinson. On the modular representation of the symmetric group.

V. Canadian Journal of Mathematics = Journal canadien de mathématiques, 7 (??):391–400, ???? 1955. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Divinsky:1955:PR

[315] Nathan Divinsky. Pseudo-regularity.

Canadian Journal of Mathematics = Journal canadien de mathématiques, 7 (??):401–410, ???? 1955. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Herstein:1955:TRC

[316] I. N. Herstein. Two remarks on the commutativity of rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 7(??):411–413, ???? 1955. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Segre:1955:OFP

[317] Beniamino Segre. Ovals in a finite projective plane. Canadian Journal of Mathematics = Journal canadien de mathématiques, 7(??):414–416, ???? 1955. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ostrom:1955:ODD

[318] T. G. Ostrom. Ovals, dualities, and Desargues’s theorem. Canadian Journal of Mathematics = Journal canadien de mathématiques, 7(??):417–431, ???? 1955. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fryer:1955:CTJ

[319] K. D. Fryer and Israel Halperin. On the coordinatization theorem of J. von Neumann. Canadian Journal of Mathematics = Journal canadien de mathématiques, 7(??):432–444, ???? 1955. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Blum:1955:SRS

[320] Richard Blum. Subspaces of Riemannian spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 7(??):445–452, ???? 1955. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Calderon:1955:FST

[321] A. P. Calderón and A. Devinatz. On Fourier–Stieltjes transforms. Canadian Journal of Mathematics = Journal canadien de mathématiques, 7(??): 453–461, ???? 1955. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hirschman:1955:WQN

[322] I. I. Hirschman, Jr. Weighted quadratic norms and Legendre polynomials. Canadian Journal of Mathematics = Journal canadien de mathématiques, 7(??): 462–482, ???? 1955. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gilmore:1955:MCR

[323] P. C. Gilmore and A. Robinson. Metamathematical considerations on the relative irreducibility of polynomials. Canadian Journal of Mathematics = Journal canadien de mathématiques, 7(??): 483–489, ???? 1955. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Higman:1955:IPM

[324] D. G. Higman. Induced and produced modules. Canadian Journal of Mathematics = Journal canadien de mathématiques, 7(??):490–508, ???? 1955. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Higman:1955:OSA

[325] D. G. Higman. On orders in separable algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 7(??):509–515, ???? 1955. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Maranda:1955:ERF

[326] J.-M. Maranda. On the equivalence of representations of finite groups by groups of automorphisms of modules over Dedekind rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 7(??):516–526, ???? 1955. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Goddard:1955:NMT

[327] L. S. Goddard. Note on a matrix theorem of A. Brauer and its extension. Canadian Journal of Mathematics = Journal canadien de mathématiques, 7 (??):527–530, ???? 1955. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Marcus:1955:SRA

[328] M. Marcus. Some results on the asymptotic behavior of linear systems. Canadian Journal of Mathematics = Journal canadien de mathématiques, 7(??): 531–538, ???? 1955. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Jaeger:1955:PDE

[329] Arno Jaeger. On partial differential equations in a field of prime characteristic. Canadian Journal of Mathematics = Journal canadien de mathématiques, 7(??):539–542, ???? 1955. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rudin:1955:CPS

[330] Mary Ellen Rudin. Countable paracompactness and Souslin’s problem. Canadian Journal of Mathematics = Journal canadien de mathématiques, 7(??): 543–547, ???? 1955. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ball:1955:NSO

[331] B. J. Ball. A note on the separability of an ordered space. Canadian Journal of Mathematics = Journal canadien de mathématiques, 7(??):548–551, ???? 1955. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Grothendieck:1955:CVM

[332] A. Grothendieck. Une caractérisation vectorielle-métrique des espaces L^{1}. (French) [A vectorial-metric characterization of L^{1} spaces]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 7(??):552–561, ???? 1955. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Koszul:1955:FHC

[333] J. L. Koszul. Sur la forme hermitienne canonique des espaces homogènes complexes. (French) [On the canonical Hermitian form of homogeneous complex spaces]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 7(??):562–576, ???? 1955. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Yoshida:1956:PER

[334] Michio Yoshida. On polynomial extensions of rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 8(??):1–2, ???? 1956. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Yoshida:1956:TZR

[335] Michio Yoshida. A theorem on Zariski rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 8(??):3–4, ???? 1956. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mills:1956:CDE

[336] W. H. Mills. Certain Diophantine equations linear in one unknown. Canadian Journal of Mathematics = Journal canadien de mathématiques, 8(??):5–12, ???? 1956. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Tutte:1956:CAG

[337] W. T. Tutte. A class of Abelian groups.

Canadian Journal of Mathematics = Journal canadien de mathématiques, 8 (??):13–28, ???? 1956. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Newman:1956:NCM

[338] Morris Newman. The normalizer of certain modular subgroups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 8(??):

29–31, ???? 1956. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wiegmann:1956:USM

[339] N. A. Wiegmann. On unitary and symmetric matrices with real quaternion elements. Canadian Journal of Mathematics = Journal canadien de mathématiques, 8(??):32–39, ???? 1956. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Goffman:1956:SRM

[340] Casper Goffman and G. M. Petersen. Submethods of regular matrix summability methods. Canadian Journal of Mathematics = Journal canadien de mathématiques, 8(??):40–46, ???? 1956. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See correction [591].

Thron:1956:ESF

[341] W. J. Thron. Entire solutions of the functional equation f(f(z)) = g(z). Canadian Journal of Mathematics = Journal canadien de mathématiques, 8 (??):47–48, ???? 1956. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rooney:1956:IFL

[342] P. G. Rooney. On an inversion formula for the Laplace transformation. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 8(??):49–52, ???? 1956. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Martin:1956:RSF

[343] A. D. Martin. A regular singular functional. Canadian Journal of Mathematics = Journal canadien de mathématiques, 8(??):53–68, ???? 1956. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Shapiro:1956:CEM

[344] J. M. Shapiro. A condition for existence of moments of infinitely divisible distributions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 8(??):69–71, ???? 1956. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hartman:1956:DDS

[345] Philip Hartman and Aurel Wintner. On disconjugate differential systems. Canadian Journal of Mathematics = Journal canadien de mathématiques, 8(??): 72–81, ???? 1956. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Diaz:1956:TR

[346] J. B. Diaz and G. S. S. Ludford. On a theorem of Le Roux. Canadian Journal of Mathematics = Journal canadien de mathématiques, 8(??):82–85, ???? 1956. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Duff:1956:UBV

[347] G. F. D. Duff. Uniqueness in boundary value problems for the second order hyperbolic equation. Canadian Journal of Mathematics = Journal canadien de mathématiques, 8(??):86–96, ???? 1956. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kazarinoff:1956:ASD

[348] N. D. Kazarinoff and R. McKelvey. Asymptotic solution of differential equations in a domain containing a regular singular point. Canadian Journal of Mathematics = Journal canadien de mathématiques, 8(??):97–104, ???? 1956. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lane:1956:DPA

[349] N. D. Lane and F. A. Sherk. Differentiable points of arcs in conformal 3-space. Canadian Journal of Mathematics = Journal canadien de mathématiques, 8(??):105–118, ???? 1956. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Schuster:1956:PPP

[350] Seymour Schuster. Pencils of polarities in projective space. Canadian Journal of Mathematics = Journal canadien de mathématiques, 8(??):119–144, ???? 1956. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kobayashi:1956:CC

[351] Shôshichi Kobayashi. On connections of Cartan. Canadian Journal of Mathematics = Journal canadien de mathématiques, 8(??):145–156, ???? 1956. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Smith:1956:GIH

[352] K. T. Smith. A generalization of an inequality of Hardy and Littlewood. Canadian Journal of Mathematics = Journal canadien de mathématiques, 8(??): 157–170, ???? 1956. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rauch:1956:HAF

[353] H. E. Rauch. Harmonic and analytic functions of several variables and the maximal theorem of Hardy and Littlewood. Canadian Journal of Mathematics = Journal canadien de mathématiques, 8(??):171–183, ???? 1956. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Jurkat:1956:EPF

[354] W. B. Jurkat. An extension problem for functions with monotonic derivatives. Canadian Journal of Mathematics = Journal canadien de mathématiques, 8(??):184–191, ???? 1956. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

McArthur:1956:RAC

[355] C. W. McArthur. On relationships amongst certain spaces of sequences in an arbitrary Banach space. Canadian Journal of Mathematics = Journal canadien de mathématiques, 8(??): 192–197, ???? 1956. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bellman:1956:QLE

[356] Richard Bellman. On a quasilinear equation. Canadian Journal of Mathematics = Journal canadien de mathématiques, 8(??):198–202, ???? 1956. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Duff:1956:MBV

[357] G. F. D. Duff. Modified boundary value problems for a quasi-linear elliptic equation. Canadian Journal of Mathematics = Journal canadien de mathématiques, 8(??):203–219, ???? 1956. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hull:1956:SAP

[358] T. E. Hull. Some algebraic properties of asymptotic power series. Canadian Journal of Mathematics = Journal canadien de mathématiques, 8(??): 220–224, ???? 1956. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Moser:1956:AE

[359] Leo Moser and Max Wyman. Asymptotic expansions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 8(??):225–233, ???? 1956. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mendelsohn:1956:ASC

[360] N. S. Mendelsohn. The asymptoic series for a certain class of permutation problems. Canadian Journal of Mathematics = Journal canadien de mathématiques, 8(??):234–244, ???? 1956. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ryser:1956:MDC

[361] H. J. Ryser. Maximal determinants in combinatorial investigations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 8(??): 245–249, ???? 1956. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kokoris:1956:CAA

[362] L. A. Kokoris. On a class of almost alternative algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 8(??):250–255, ???? 1956. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

deGroot:1956:OIR

[363] J. de Groot. Orthogonal isomorphic representations of free groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 8(??): 256–262, ???? 1956. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Tobin:1956:TBH

[364] Seán Tobin. On a theorem of Baer and Higman. Canadian Journal of Mathematics = Journal canadien de mathématiques, 8(??):263–270, ???? 1956. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Curtis:1956:CRE

[365] C. W. Curtis. On commuting rings of endomorphisms. Canadian Journal of Mathematics = Journal canadien de mathématiques, 8(??):271–292, ???? 1956. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Barnes:1956:CSS

[366] E. S. Barnes. The covering of space by spheres. Canadian Journal of Mathematics = Journal canadien de mathématiques, 8(??):293–304, ???? 1956. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Touchard:1956:NEN

[367] Jacques Touchard. Nombres exponentiels et nombres de Bernoulli. (French) [Exponential numbers and Bernoulli numbers]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 8(??):305–320, ???? 1956. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wyman:1956:SPT

[368] Max Wyman and Leo Moser. On some polynomials of Touchard. Canadian Journal of Mathematics = Journal canadien de mathématiques, 8(??): 321–322, ???? 1956. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Yaqub:1956:TRL

[369] Adil Yaqub. On the theory of ring-logics. Canadian Journal of Mathematics = Journal canadien de mathématiques, 8 (??):323–328, ???? 1956. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Reiner:1956:MMD

[370] Irving Reiner. Maschke modules over Dedekind rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 8(??):329–334, ???? 1956. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kleinfeld:1956:SAR

[371] Erwin Kleinfeld. Standard and accessible rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 8(??):335–340, ???? 1956. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Drazin:1956:ADR

[372] M. P. Drazin. Algebraic and diagonable rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 8 (??):341–354, ???? 1956. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Amitsur:1956:RPR

[373] S. A. Amitsur. Radicals of polynomial rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 8 (??):355–361, ???? 1956. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Edge:1956:COP

[374] W. L. Edge. Conics and orthogonal projectivities in a finite plane. Canadian Journal of Mathematics = Journal canadien de mathématiques, 8(??): 362–382, ???? 1956. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Derry:1956:CHS

[375] Douglas Derry. Convex hulls of simple space curves. Canadian Journal of Mathematics = Journal canadien de mathématiques, 8(??):383–388, ???? 1956. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Handest:1956:CHG

[376] Frans Handest. Constructions in hyperbolic geometry. Canadian Journal of Mathematics = Journal canadien de mathématiques, 8(??):389–394, ???? 1956. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Banaschewski:1956:LCE

[377] Bernhard Banaschewski. Local connectedness of extension spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 8(??): 395–398, ???? 1956. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ford:1956:MFT

[378] L. R. Ford, Jr. and D. R. Fulkerson. Maximal flow through a network. Canadian Journal of Mathematics = Journal canadien de mathématiques, 8(??): 399–404, ???? 1956. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gilbert:1956:ELG

[379] E. N. Gilbert. Enumeration of labelled graphs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 8(??):405–411, ???? 1956. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dulmage:1956:CIZ

[380] A. L. Dulmage and J. E. L. Peck. Certain infinite zero-sum two-person games. Canadian Journal of Mathematics = Journal canadien de mathématiques, 8 (??):412–416, ???? 1956. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ellis:1956:BPV

[381] H. W. Ellis. On the basis problem for vector valued function spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 8(??): 417–422, ???? 1956. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Jenkins:1956:EBL

[382] J. A. Jenkins. On explicit bounds in Landau’s theorem. Canadian Journal of Mathematics = Journal canadien de mathématiques, 8(??):423–425, ????

1956. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Miles:1956:CPL

[383] E. P. Miles, Jr. and Ernest Williams.

The Cauchy problem for linear partial differential equations with restricted boundary conditions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 8(??):426–431, ???? 1956. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ericksen:1956:IHC

[384] J. L. Ericksen and R. A. Toupin.

Implications of Hadamard’s conditions for elastic stability with respect to uniqueness theorems. Canadian Journal of Mathematics = Journal canadien de mathématiques, 8(??):432–436, ???? 1956. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sumner:1956:GAO

[385] D. B. Sumner. A generalized averaging operator. Canadian Journal of Mathematics = Journal canadien de mathématiques, 8(??):437–446, ???? 1956. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Porcelli:1956:NST

[386] Pasquale Porcelli. Note on a Stieltjes type of inversion. Canadian Journal of Mathematics = Journal canadien de mathématiques, 8(??):447–448, ???? 1956. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gonshor:1956:STC

[387] Harry Gonshor. Spectral theory for a class of nonnormal operators. Canadian Journal of Mathematics = Journal canadien de mathématiques, 8(??): 449–461, ???? 1956. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dye:1956:GPO

[388] H. A. Dye and R. S. Phillips. Groups of positive operators. Canadian Journal of Mathematics = Journal canadien de mathématiques, 8(??):462–486, ???? 1956. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rund:1956:HFS

[389] Hanno Rund. Hypersurfaces of a Finsler space. Canadian Journal of Mathematics = Journal canadien de mathématiques, 8(??):487–503, ???? 1956. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Beesack:1956:ZSS

[390] P. R. Beesack and Binyamin Schwarz. On the zeros of solutions of secondorder linear differential equations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 8(??): 504–515, ???? 1956. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Leblanc:1956:EMM

[391] L. Leblanc and G. E. Fox. On the extension of measure by the method of Borel. Canadian Journal of Mathematics = Journal canadien de mathématiques, 8 (??):516–523, ???? 1956. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Marcus:1956:EPH

[392] M. Marcus and J. L. McGregor. Extremal properties of Hermitian matrices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 8(??):524–531, ???? 1956. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mendelsohn:1956:NDP

[393] N. S. Mendelsohn. Non-Desarguesian projective plane geometries which satisfy the harmonic point axiom. Canadian Journal of Mathematics = Journal canadien de mathématiques, 8(??): 532–562, ???? 1956. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ostrom:1956:DTF

[394] T. G. Ostrom. Double transitivity in finite projective planes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 8(??):563–567, ???? 1956. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Carlitz:1956:RCL

[395] L. Carlitz. Resolvents of certain linear groups in a finite field. Canadian Journal of Mathematics = Journal canadien de mathématiques, 8(??):568–579, ???? 1956. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Steinberg:1956:PPR

[396] Robert Steinberg. Prime power representations of finite linear groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 8(??): 580–591, ???? 1956. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Jones:1956:ITQ

[397] B. W. Jones and G. L. Watson. On indefinite ternary quadratic forms. Canadian Journal of Mathematics = Journal canadien de mathématiques, 8(??): 592–608, ???? 1956. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mirkil:1957:NCP

[398] H. Mirkil. New characterizations of polyhedral cones. Canadian Journal of Mathematics = Journal canadien de mathématiques, 9(??):1–4, ???? 1957. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hsiung:1957:SGT

[399] Chuan-Chih Hsiung. Some global theorems on hyper-surfaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 9(??):5–14, ???? 1957. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Paige:1957:NMG

[400] Lowell J. Paige. A note on the Mathieu groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 9(??):15–18, ???? 1957. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Higman:1957:RC

[401] D. G. Higman. Relative cohomology. Canadian Journal of Mathematics = Journal canadien de mathématiques, 9 (??):19–34, ???? 1957. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Yoshida:1957:SRN

[402] Michio Yoshida. Some remarks on Noetherian rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 9(??):35–37, ???? 1957. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Banaschewski:1957:SDZ

[403] Bernhard Banaschewski. Spaces of dimension zero. Canadian Journal of Mathematics = Journal canadien de mathématiques, 9(??):38–46, ???? 1957. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Butson:1957:MEB

[404] A. T. Butson. Matrices with elements in a Boolean ring. Canadian Journal of Mathematics = Journal canadien de mathématiques, 9(??):47–59, ???? 1957. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Schneider:1957:CP

[405] Hans Schneider. Characteristic polynomials. Canadian Journal of Mathematics = Journal canadien de mathématiques, 9(??):60–67, ???? 1957. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Newman:1957:STA

[406] Morris Newman. Some theorems about psbr(n). Canadian Journal of Mathematics = Journal canadien de mathématiques, 9(??):68–70, ???? 1957. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Newman:1957:CPD

[407] Morris Newman and Olga Taussky. Classes of positive definite unimodular circulants. Canadian Journal of Mathematics = Journal canadien de mathématiques, 9(??):71–73, ???? 1957. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cohen:1957:SPL

[408] Eckford Cohen. Simultaneous pairs of linear and quadratic equations in a Galois field. Canadian Journal of Mathematics = Journal canadien de mathématiques, 9(??):74–78, ???? 1957. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Jerison:1957:SAG

[409] Meyer Jerison. The set of all generalized limits of bounded sequences. Canadian Journal of Mathematics = Journal canadien de mathématiques, 9 (??):79–89, ???? 1957. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Davis:1957:NAS

[410] Harry F. Davis. A note on asymptotic series. Canadian Journal of Mathematics = Journal canadien de mathématiques, 9(??):90–95, ???? 1957. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Henstock:1957:WPS

[411] Ralph Henstock. On Ward’s Perron– Stieltjes integral. Canadian Journal of Mathematics = Journal canadien de mathématiques, 9(??):96–109, ???? 1957. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fox:1957:GCP

[412] Charles Fox. A generalization of the Cauchy principal value. Canadian Journal of Mathematics = Journal canadien de mathématiques, 9(??):110–117, ???? 1957. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kreyszig:1957:ZFI

[413] Erwin Kreyszig. On the zeros of the Fresnel integrals. Canadian Journal of Mathematics = Journal canadien de mathématiques, 9(??):118–131, ???? 1957. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Carter:1957:MMP

[414] D. S. Carter. A minimum-maximum problem for differential expressions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 9 (??):132–140, ???? 1957. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Duff:1957:MPN

[415] G. F. D. Duff. A mixed problem for normal hyperbolic linear partial differential equations of second order. Canadian Journal of Mathematics = Journal canadien de mathématiques, 9(??): 141–160, ???? 1957. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Duff:1957:HMP

[416] G. F. D. Duff. Hyperbolic mixed problems for harmonic tensors. Canadian Journal of Mathematics = Journal canadien de mathématiques, 9(??): 161–179, ???? 1957. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Brafman:1957:SGF

[417] Fred Brafman. Some generating functions for Laguerre and Hermite polynomials. Canadian Journal of Mathematics = Journal canadien de mathématiques, 9(??):180–187, ???? 1957. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Carlitz:1957:SPT

[418] Leonard Carlitz. Some polynomials of Touchard connected with the Bernoulli numbers. Canadian Journal of Mathematics = Journal canadien de mathématiques, 9(??):188–190, ???? 1957. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Brafman:1957:TP

[419] Fred Brafman. On Touchard polynomials. Canadian Journal of Mathematics = Journal canadien de mathématiques, 9(??):191–193, ???? 1957. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Moser:1957:AEI

[420] Leo Moser and Max Wyman. Asymptotic expansions. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 9(??):194–209, ???? 1957. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ford:1957:SAF

[421] L. R. Ford, Jr. and D. R. Fulkerson. A simple algorithm for finding maximal network flows and an application to the Hitchcock problem. Canadian Journal of Mathematics = Journal canadien de mathématiques, 9(??):210–218, ???? 1957. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Roberts:1957:DP

[422] J. B. Roberts. On a Diophantine problem. Canadian Journal of Mathematics = Journal canadien de mathématiques, 9(??):219–222, ???? 1957. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Petersen:1957:SSS

[423] G. M. Petersen. Sets and sub-series.

Canadian Journal of Mathematics = Journal canadien de mathématiques, 9 (??):223–224, ???? 1957. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dirac:1957:SPM

[424] G. A. Dirac. Short proof of a mapcolour theorem. Canadian Journal of Mathematics = Journal canadien de mathématiques, 9(??):225–226, ???? 1957. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Carter:1957:CMM

[425] D. S. Carter. Correction to a minimum-maximum problem for differential expressions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 9(??):226, ???? 1957. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Frucht:1957:ULB

[426] Robert Frucht. Upper and lower bounds for the area of a triangle for whose sides two symmetric functions are known. Canadian Journal of Mathematics = Journal canadien de mathématiques, 9 (??):227–231, ???? 1957. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dempster:1957:MDT

[427] A. P. Dempster. The minimum of a definite ternary quadratic form. Canadian Journal of Mathematics = Journal canadien de mathématiques, 9(??): 232–234, ???? 1957. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Barnes:1957:PES

[428] E. S. Barnes. The perfect and extreme senary forms. Canadian Journal of Mathematics = Journal canadien de mathématiques, 9(??):235–242, ???? 1957. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Coxeter:1957:GGU

[429] H. S. M. Coxeter. Groups generated by unitary reflections of period two. Canadian Journal of Mathematics = Journal canadien de mathématiques, 9(??): 243–272, ???? 1957. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Shephard:1957:ADR

[430] G. C. Shephard. Abstract definitions for reflection groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 9(??):273–276, ???? 1957. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rankin:1957:DAH

[431] R. A. Rankin. Diophantine approximation and horocyclic groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 9(??): 277–290, ???? 1957. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

deGroot:1957:EAG

[432] J. de Groot. Equivalent abelian groups.

Canadian Journal of Mathematics = Journal canadien de mathématiques, 9 (??):291–297, ???? 1957. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fan:1957:ICH

[433] Ky Fan and Gordon Pall. Imbedding conditions for Hermitian and normal matrices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 9(??):298–304, ???? 1957. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Marcus:1957:ISF

[434] M. Marcus and L. Lopes. Inequalities for symmetric functions and Hermitian matrices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 9(??):305–312, ???? 1957. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Marcus:1957:MPK

[435] M. Marcus and B. N. Moyls. On the maximum principle of Ky fan. Canadian Journal of Mathematics = Journal canadien de mathématiques, 9(??): 313–320, ???? 1957. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sade:1957:GAP

[436] A. Sade. Groupoïdes automorphes par le groupe cyclique. (French) [Automorphic groupoids by the cyclic group]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 9(??):

321–335, ???? 1957. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wiegmann:1957:SGB

[437] N. A. Wiegmann. Some generalizations of Burnside’s theorem. Canadian Journal of Mathematics = Journal canadien de mathématiques, 9(??):336–346, ???? 1957. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Steinberg:1957:PPR

[438] Robert Steinberg. Prime power representations of finite linear groups. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 9 (??):347–351, ???? 1957. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

McCoy:1957:CCI

[439] N. H. McCoy. Certain classes of ideals in polynomial rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 9(??):352–362, ???? 1957. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Roberts:1957:BCR

[440] J. B. Roberts. On binomial coefficient residues. Canadian Journal of Mathematics = Journal canadien de mathématiques, 9(??):363–370, ???? 1957. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ryser:1957:CPM

[441] H. J. Ryser. Combinatorial properties of matrices of zeros and ones. Canadian Journal of Mathematics = Journal canadien de mathématiques, 9(??):

371–377, ???? 1957. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hughes:1957:CND

[442] D. R. Hughes. A class of non-Desarguesian projective planes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 9(??): 378–388, ???? 1957. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ostrom:1957:TPP

[443] T. G. Ostrom. Transitivities in projective planes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 9(??):389–399, ???? 1957. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See correction [515].

Wolk:1957:DCF

[444] E. S. Wolk. Dedekind completeness and a fixed-point theorem. Canadian Journal of Mathematics = Journal canadien de mathématiques, 9(??):400–405, ???? 1957. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Aissen:1957:COS

[445] M. I. Aissen. Cyclically ordered sets.

Canadian Journal of Mathematics = Journal canadien de mathématiques, 9 (??):406–412, ???? 1957. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Price:1957:CGO

[446] J. J. Price. Certain groups of orthonormal step functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 9(??):413–425, ????

1957. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rudin:1957:CIA

[447] Walter Rudin. The closed ideals in an algebra of analytic functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 9(??): 426–434, ???? 1957. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Schatz:1957:RBA

[448] J. A. Schatz. Representation of Banach algebras with an involution. Canadian Journal of Mathematics = Journal canadien de mathématiques, 9(??): 435–442, ???? 1957. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mansfield:1957:CPN

[449] M. J. Mansfield. On countably paracompact normal spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 9(??):443–449, ???? 1957. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Peck:1957:GCS

[450] J. E. L. Peck and A. L. Dulmage. Games on a compact set. Canadian Journal of Mathematics = Journal canadien de mathématiques, 9(??):450–458, ???? 1957. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rooney:1957:IGT

[451] P. G. Rooney. On the inversion of the Gauss transformation. Canadian Journal of Mathematics = Journal canadien de mathématiques, 9(??):459–464, ????

1957. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gordon:1957:PLF

[452] H. Gordon and E. R. Lorch. The projection of a linear functional on the manifold of integrals. Canadian Journal of Mathematics = Journal canadien de mathématiques, 9(??):465–474, ???? 1957. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Morse:1957:PML

[453] Marston Morse and William Transue.

Products of a C-measure and a locally integrable mapping. Canadian Journal of Mathematics = Journal canadien de mathématiques, 9(??):475–486, ???? 1957. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rajagopal:1957:TTR

[454] C. T. Rajagopal. A Tauberian theorem for the Riemann–Liouville integral of integer order. Canadian Journal of Mathematics = Journal canadien de mathématiques, 9(??):487–499, ???? 1957. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kreyszig:1957:CFF

[455] Erwin Kreyszig. On the complementary functions of the Fresnel integrals. Canadian Journal of Mathematics = Journal canadien de mathématiques, 9(??): 500–510, ???? 1957. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

deGroot:1957:MCD

[456] J. de Groot. On a metric that characterizes dimension. Canadian Journal of Mathematics = Journal canadien de mathématiques, 9(??):511–514, ???? 1957. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sabidussi:1957:GGG

[457] Gert Sabidussi. Graphs with given group and given graph-theoretical properties. Canadian Journal of Mathematics = Journal canadien de mathématiques, 9 (??):515–525, ???? 1957. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Watson:1957:EQF

[458] G. L. Watson. The equivalence of quadratic forms. Canadian Journal of Mathematics = Journal canadien de mathématiques, 9(??):526–548, ???? 1957. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Newman:1957:CCM

[459] Morris Newman. Congruences for the coefficients of modular forms and some new congruences for the partition function. Canadian Journal of Mathematics = Journal canadien de mathématiques, 9(??):549–552, ???? 1957. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Pearl:1957:CP

[460] M. H. Pearl. On Cayley’s parameterization. Canadian Journal of Mathematics = Journal canadien de mathématiques, 9(??):553–562, ???? 1957. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Isbell:1957:SRC

[461] J. R. Isbell. Some remarks concerning categories and subspaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 9(??): 563–577, ???? 1957. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ward:1957:CSL

[462] L. E. Ward, Jr. Completeness in semi-lattices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 9(??):578–582, ???? 1957. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Herstein:1957:CCR

[463] I. N. Herstein. A condition for the commutativity of rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 9(??):583–586, ???? 1957. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Feit:1957:SFG

[464] Walter Feit. On the structure of Frobenius groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 9(??):587–596, ???? 1957. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Maranda:1957:FR

[465] J.-M. Maranda. Factorization rings.

Canadian Journal of Mathematics = Journal canadien de mathématiques, 9 (??):597–623, ???? 1957. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Littlewood:1958:IPF

[466] D. E. Littlewood. The inner plethysm of S-functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 10(??):1–16, ???? 1958.

CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Littlewood:1958:PPC

[467] D. E. Littlewood. Products and plethysms of characters with orthogonal, symplectic and symmetric groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 10 (??):17–32, ???? 1958. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Davis:1958:SFL

[468] Robert L. Davis. A special formula for the Lie character. Canadian Journal of Mathematics = Journal canadien de mathématiques, 10(??):33–38, ???? 1958. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Jans:1958:RTA

[469] J. P. Jans. The representation type of algebras and subalgebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 10(??):39–44, ???? 1958. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Lambek:1958:GTZ

[470] Joachim Lambek. Goursat’s theorem and the Zassenhaus lemma. Canadian Journal of Mathematics = Journal canadien de mathématiques, 10(??): 45–56, ???? 1958. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ryser:1958:TRM

[471] H. J. Ryser. The term rank of a matrix. Canadian Journal of Mathematics = Journal canadien de mathématiques, 10(??):57–65, ???? 1958. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Vartak:1958:HMI

[472] Manohar N. Vartak. On the Hasse– Minkowski invariant of the Kronecker product of matrices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 10(??):66–72, ???? 1958. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Stanton:1958:FDS

[473] R. G. Stanton and D. A. Sprott. A family of difference sets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 10(??):73–77, ???? 1958. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ford:1958:NFS

[474] L. R. Ford, Jr. and D. R. Fulkerson.

Network flow and systems of representatives. Canadian Journal of Mathematics = Journal canadien de mathématiques, 10(??):78–84, ???? 1958. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Handscomb:1958:RDT

[475] D. C. Handscomb. On the random disorientation of two cubes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 10(??): 85–88, ???? 1958. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Helfenstein:1958:GGM

[476] H. G. Helfenstein. Geodesic groups of minimal surfaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 10(??):89–96, ???? 1958. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gonshor:1958:STC

[477] Harry Gonshor. Spectral theory for a class of nonnormal operators. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 10(??): 97–102, ???? 1958. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Schmidt:1958:CMV

[478] Wolfgang Schmidt. On the convergence of mean values over lattices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 10(??): 103–110, ???? 1958. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Keogh:1958:GTT

[479] F. R. Keogh and G. M. Petersen. A generalized Tauberian theorem. Canadian Journal of Mathematics = Journal canadien de mathématiques, 10(??): 111–114, ???? 1958. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kober:1958:EBI

[480] H. Kober. On the existence of the Burkill integral. Canadian Journal of Mathematics = Journal canadien de mathématiques, 10(??):115–121, ???? 1958. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Boas:1958:GAO

[481] R. P. Boas, Jr. On generalized averaging operators. Canadian Journal of Mathematics = Journal canadien de mathématiques, 10(??):122–126, ???? 1958. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Duff:1958:MPL

[482] G. F. D. Duff. Mixed problems for linear systems of first order equations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 10(??): 127–160, ???? 1958. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Robertson:1958:SDS

[483] M. S. Robertson. Schlicht Dirichlet series. Canadian Journal of Mathematics = Journal canadien de mathématiques, 10(??):161–176, ???? 1958. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rooney:1958:LTG

[484] P. G. Rooney. Laplace transforms and generalized Laguerre polynomials. Canadian Journal of Mathematics = Journal canadien de mathématiques, 10 (??):177–182, ???? 1958. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kreyszig:1958:SRB

[485] Erwin Kreyszig. On some relations between partial and ordinary differential equations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 10(??):183–190, ???? 1958. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Roberts:1958:NPT

[486] J. B. Roberts. A new proof of a theorem of Lehmer. Canadian Journal of Mathematics = Journal canadien de mathématiques, 10(??):191–194, ???? 1958. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Shukla:1958:CTN

[487] H. S. Shukla. Certain transformations of nearly-poised bilateral hypergeometric series of special type. Canadian Journal of Mathematics = Journal canadien de mathématiques, 10(??):195–201, ???? 1958. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Golomb:1958:CFC

[488] S. W. Golomb, Basil Gordon, and L. R. Welch. Comma-free codes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 10(??): 202–209, ???? 1958. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kelly:1958:NOL

[489] L. M. Kelly and W. O. J. Moser. On the number of ordinary lines determined by n points. Canadian Journal of Mathematics = Journal canadien de mathématiques, 10(??):210–219, ???? 1958. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Steinberg:1958:NSP

[490] Robert Steinberg. On the number of sides of a Petrie polygon. Canadian Journal of Mathematics = Journal canadien de mathématiques, 10(??): 220–221, ???? 1958. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Blum:1958:CLT

[491] J. R. Blum, H. Chernoff, M. Rosenblatt, and H. Teicher. Central limit theorems for interchangeable processes. Canadian Journal of Mathematics = Journal canadien de mathématiques, 10(??): 222–229, ???? 1958. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mendelsohn:1958:SGP

[492] N. S. Mendelsohn and A. L. Dulmage.

Some generalizations of the problem of distinct representatives. Canadian Journal of Mathematics = Journal canadien de mathématiques, 10(??):230–241, ???? 1958. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wallace:1958:AAC

[493] A. H. Wallace. Algebraic approximation of curves. Canadian Journal of Mathematics = Journal canadien de mathématiques, 10(??):242–278, ???? 1958. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Chang:1958:CPM

[494] Bomshik Chang, S. A. Jennings, and Rimhak Ree. On certain pairs of matrices which generate free groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 10(??): 279–284, ???? 1958. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dubois:1958:NDA

[495] D. W. Dubois and A. Steger. A note on division algorithms in imaginary quadratic number fields. Canadian Journal of Mathematics = Journal canadien de mathématiques, 10(??): 285–286, ???? 1958. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Auslander:1958:DIP

[496] Maurice Auslander and Alex Rosenberg. Dimension of ideals in polynomial rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 10 (??):287–293, ???? 1958. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sade:1958:GAP

[497] A. Sade. Groupoïdes automorphes par le groupe géométrique et quasigroupes “endo”. (French) [Automorphic groupoids from the geometric group and “endo” quasigroups]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 10(??):294–320, ???? 1958. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dixmier:1958:RUG

[498] Jacques Dixmier. Sur les représentations unitaries des groupes de Lie nilpotents. III. (French) [On the unitary representations of nilpotent Lie groups. III]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 10(??): 321–348, ???? 1958. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Coleman:1958:BNS

[499] A. J. Coleman. The Betti numbers of the simple Lie groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 10(??):349–356, ???? 1958. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dekker:1958:RS

[500] J. C. E. Dekker and J. Myhill. Retraceable sets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 10(??):357–373, ???? 1958. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Faith:1958:CDR

[501] Carl C. Faith. On conjugates in division rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 10(??):374–380, ???? 1958. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ellis:1958:CS

[502] H. W. Ellis. On the MT^{∗}- and λconjugates of L^{λ} spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 10(??):381–391, ???? 1958. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gosselin:1958:CTD

[503] Richard P. Gosselin. A convergence theorem for double L^{2} Fourier series. Canadian Journal of Mathematics = Journal canadien de mathématiques, 10(??): 392–398, ???? 1958. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Snow:1958:IVV

[504] D. O. Snow. On integration of vectorvalued functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 10(??):399–412, ???? 1958. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kassimatis:1958:FWG

[505] C. Kassimatis. Functions which have generalized Riemann derivatives. Canadian Journal of Mathematics = Journal canadien de mathématiques, 10(??):

413–420, ???? 1958. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rooney:1958:STD

[506] P. G. Rooney. On some theorems of Doetsch. Canadian Journal of Mathematics = Journal canadien de mathématiques, 10(??):421–430, ???? 1958. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Brauer:1958:STD

[507] Fred Brauer. Spectral theory for the differential equation Lu = λMu. Canadian Journal of Mathematics = Journal canadien de mathématiques, 10(??): 431–446, ???? 1958. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kemp:1958:SBV

[508] R. R. D. Kemp. A singular boundary value problem for a non-self-adjoint differential operator. Canadian Journal of Mathematics = Journal canadien de mathématiques, 10(??):447–462, ???? 1958. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Henrici:1958:PTK

[509] Peter Henrici. On the product of two Kummer series. Canadian Journal of Mathematics = Journal canadien de mathématiques, 10(??):463–467, ???? 1958. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wyman:1958:PM

[510] Max Wyman and Leo Moser. On the problème des ménages. Canadian Journal of Mathematics = Journal canadien de mathématiques, 10(??):468–480, ????

1958. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Tutte:1958:CNR

[511] W. T. Tutte. The chords of the non-ruled quadric in P G(3, 3). Canadian Journal of Mathematics = Journal canadien de mathématiques, 10(??): 481–483, ???? 1958. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Coxeter:1958:CNR

[512] H. S. M. Coxeter. The chords of the non-ruled quadric in P G(3, 3). Canadian Journal of Mathematics = Journal canadien de mathématiques, 10(??): 484–488, ???? 1958. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mandan:1958:CEE

[513] Sahib Ram Mandan. An S-configuration in Euclidean and elliptic n-space. Canadian Journal of Mathematics = Journal canadien de mathématiques, 10(??): 489–501, ???? 1958. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Bottema:1958:MTH

[514] O. Bottema. On the medians of a triangle in hyperbolic geometry. Canadian Journal of Mathematics = Journal canadien de mathématiques, 10(??): 502–506, ???? 1958. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ostrom:1958:CTP

[515] T. G. Ostrom. Correction to “Transitivities in projective planes”. Canadian Journal of Mathematics = Journal canadien de mathématiques, 10(??): 507–512, ???? 1958. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See [443].

Harary:1958:NDG

[516] Frank Harary. On the number of dissimilar graphs between a given graphsubgraph pair. Canadian Journal of Mathematics = Journal canadien de mathématiques, 10(??):513–516, ???? 1958. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dulmage:1958:CBG

[517] A. L. Dulmage and N. S. Mendelsohn. Coverings of bipartite graphs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 10(??): 517–534, ???? 1958. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Englefield:1958:OFS

[518] M. J. Englefield. Oriented flat submanifolds in an affine space. Canadian Journal of Mathematics = Journal canadien de mathématiques, 10(??):535–546, ???? 1958. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hartmanis:1958:LT

[519] Juris Hartmanis. On the lattice of topologies. Canadian Journal of Mathematics = Journal canadien de mathématiques, 10(??):547–553, ???? 1958. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rothberger:1958:EEE

[520] F. Rothberger. Exemple effectif d’un ensemble transfiniment non-projectif.

(French) [Practical example of a transfinitely non-projective set]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 10(??): 554–560, ???? 1958. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cohen:1958:CRA

[521] Eckford Cohen. Congruence representations in algebraic number fields. II. simultaneous linear and quadratic congruences. Canadian Journal of Mathematics = Journal canadien de mathématiques, 10(??):561–571, ???? 1958. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Keogh:1958:WDS

[522] F. R. Keogh, B. Lawton, and G. M.

Petersen. Well distributed sequences. Canadian Journal of Mathematics = Journal canadien de mathématiques, 10 (??):572–576, ???? 1958. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Newman:1958:FIC

[523] Morris Newman. Further identities and congruences for the coefficients of modular forms. Canadian Journal of Mathematics = Journal canadien de mathématiques, 10(??):577–586, ???? 1958. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Putnam:1958:BMN

[524] C. R. Putnam. On bounded matrices with non-negative elements. Canadian Journal of Mathematics = Journal canadien de mathématiques, 10(??): 587–591, ???? 1958. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

vanderSluis:1958:OPH

[525] A. van der Sluis. Orthogonal polynomials and hypergeometric series. Canadian Journal of Mathematics = Journal canadien de mathématiques, 10(??): 592–612, ???? 1958. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rooney:1958:IGT

[526] P. G. Rooney. On the inversion of the Gauss transformation. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 10(??): 613–616, ???? 1958. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Jeffery:1958:GIR

[527] R. L. Jeffery. Generalized integrals with respect to functions of bounded variation. Canadian Journal of Mathematics = Journal canadien de mathématiques, 10(??):617–626, ???? 1958. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Jenkins:1958:DC

[528] James A. Jenkins. On the Denjoy conjecture. Canadian Journal of Mathematics = Journal canadien de mathématiques, 10(??):627–631, ???? 1958. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

MacCamy:1958:BP

[529] R. C. MacCamy. On Babinet’s principle. Canadian Journal of Mathematics = Journal canadien de mathématiques, 10(??):632–640, ???? 1958. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wallace:1959:AEA

[530] Andrew H. Wallace. Analytic equivalence of algebroid curves. Canadian Journal of Mathematics = Journal canadien de mathématiques, 11(??): 1–17, ???? 1959. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Horadam:1959:IAB

[531] A. F. Horadam. Involutions associated with the Burkhardt configuration in [4]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 11(??):18–33, ???? 1959. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Erdos:1959:GTP

[532] P. Erdős. Graph theory and probability. Canadian Journal of Mathematics = Journal canadien de mathématiques, 11(??):34–38, ???? 1959. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gorenstein:1959:CFG

[533] Daniel Gorenstein. A class of Frobenius groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 11(??):39–47, ???? 1959. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Pearl:1959:FEC

[534] Martin Pearl. A further extension of Cayley’s parametrization. Canadian Journal of Mathematics = Journal canadien de mathématiques, 11(??):

48–50, ???? 1959. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Zemmer:1959:SDA

[535] Joseph L. Zemmer. Some G division algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 11(??):51–58, ???? 1959. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Nagao:1959:ROR

[536] Hirosi Nagao. A remark on the orthogonality relations in the representation theory of finite groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 11(??):59–60, ???? 1959. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Marcus:1959:LTAa

[537] Marvin Marcus and B. N. Moyls. Linear transformations on algebras of matrices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 11(??):61–66, ???? 1959. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dekker:1959:FPC

[538] Th. J. Dekker. On free products of cyclic rotation groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 11(??):67–69, ???? 1959. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Warner:1959:CCP

[539] Seth Warner. Characters of Cartesian products of algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 11(??):70–79, ????

1959. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Brainerd:1959:RCF

[540] Barron Brainerd. F -rings of continuous functions. I. Canadian Journal of Mathematics = Journal canadien de mathématiques, 11(??):80–86, ???? 1959. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Banaschewski:1959:CEF

[541] Bernhard Banaschewski. On certain extensions of function rings. Canadian Journal of Mathematics = Journal canadien de mathématiques, 11(??): 87–96, ???? 1959. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hartmanis:1959:LTG

[542] Juris Hartmanis. Lattice theory of generalized partitions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 11(??):97–106, ???? 1959. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Krule:1959:CBR

[543] I. S. Krule. Concerning binary relations on connected ordered spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 11(??): 107–111, ???? 1959. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Tulcea:1959:COO

[544] C. Ionescu Tulcea. On a class of operators occurring in the theory of chains of infinite order. Canadian Journal of Mathematics = Journal canadien de mathématiques, 11(??):112–121, ????

1959. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Brown:1959:TRF

[545] Richard K. Brown. Typically-real functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 11(??):122–130, ???? 1959. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dawson:1959:CFA

[546] David F. Dawson. Continued fractions with absolutely convergent even or odd part. Canadian Journal of Mathematics = Journal canadien de mathématiques, 11(??):131–140, ???? 1959. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Weisner:1959:GFH

[547] Louis Weisner. Generating functions for Hermite functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 11(??):141–147, ???? 1959. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Weisner:1959:GFB

[548] Louis Weisner. Generating functions for Bessel functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 11(??):148–155, ???? 1959. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ragab:1959:SPB

[549] F. M. Ragab. Series of products of Bessel polynomials. Canadian Journal of Mathematics = Journal canadien de mathématiques, 11(??):156–160, ???? 1959. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Erdos:1959:SRP

[550] P. Erdős. Some remarks on prime factors of integers. Canadian Journal of Mathematics = Journal canadien de mathématiques, 11(??):161–167, ???? 1959. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rooney:1959:RFF

[551] P. G. Rooney. On the representation of functions as Fourier transforms. Canadian Journal of Mathematics = Journal canadien de mathématiques, 11(??): 168–174, ???? 1959. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Neville:1959:EIT

[552] E. H. Neville. The elliptic integrals of the third kind. Canadian Journal of Mathematics = Journal canadien de mathématiques, 11(??):175–194, ???? 1959. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Duff:1959:MPH

[553] G. F. D. Duff. Mixed problems for hyperbolic equations of general order. Canadian Journal of Mathematics = Journal canadien de mathématiques, 11 (??):195–221, ???? 1959. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Cohn:1959:IRC

[554] Richard M. Cohn. An improved result concerning singular manifolds of difference polynomials. Canadian Journal of Mathematics = Journal canadien de mathématiques, 11(??):222–234, ???? 1959. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Eliopoulos:1959:SGM

[555] H. A. Eliopoulos. Subspaces of a generalized metric space. Canadian Journal of Mathematics = Journal canadien de mathématiques, 11(??):235–255, ???? 1959. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Woods:1959:ICB

[556] A. C. Woods. On the irreducibilty of convex bodies. Canadian Journal of Mathematics = Journal canadien de mathématiques, 11(??):256–261, ???? 1959. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Nettleton:1959:DSC

[557] R. E. Nettleton, K. Goldberg, and M. S. Green. Dense subgraphs and connectivity. Canadian Journal of Mathematics = Journal canadien de mathématiques, 11 (??):262–268, ???? 1959. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dulmage:1959:TSR

[558] A. L. Dulmage and N. S. Mendelsohn. The term and stochastic ranks of a matrix. Canadian Journal of Mathematics = Journal canadien de mathématiques, 11(??):269–279, ???? 1959. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Higgins:1959:DTS

[559] P. J. Higgins. Disjoint transversals of subsets. Canadian Journal of Mathematics = Journal canadien de mathématiques, 11(??):280–285, ???? 1959. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Leader:1959:SAT

[560] Solomon Leader. Separation and approximation in topological vector lattices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 11(??):286–296, ???? 1959. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gelbaum:1959:TPB

[561] Bernard R. Gelbaum. Tensor products of Banach algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 11(??):297–310, ???? 1959. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Gorenstein:1959:CSG

[562] Daniel Gorenstein and I. N. Herstein. A class of solvable groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 11(??):311–320, ???? 1959. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dixmier:1959:RUG

[563] Jacques Dixmier. Sur les représentations unitaires des groupes de Lie nilpotents. IV. (French) [On the unitary representations of nilpotent Lie groups. IV]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 11(??): 321–344, ???? 1959. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ree:1959:NGW

[564] Rimhak Ree. Note on generalized Witt algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 11(??):345–352, ???? 1959. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Baer:1959:SI

[565] Reinhold Baer. Supersoluble immersion. Canadian Journal of Mathematics = Journal canadien de mathématiques, 11(??):353–369, ???? 1959. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Walker:1959:NLR

[566] Elbert A. Walker. Non-linear recursive sequences. Canadian Journal of Mathematics = Journal canadien de mathématiques, 11(??):370–378, ???? 1959. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Marcus:1959:EPH

[567] M. Marcus, B. N. Moyls, and R. Westwick. Extremal properties of Hermitian matrices. II. Canadian Journal of Mathematics = Journal canadien de mathématiques, 11(??):379–382, ???? 1959. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Marcus:1959:LTAb

[568] Marvin Marcus and Roger Purves. Linear transformations on algebras of matrices: the invariance of the elementary symmetric functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 11(??):383–396, ???? 1959. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Heider:1959:PDI

[569] L. J. Heider. Prime dual ideals in Boolean algebras. Canadian Journal of Mathematics = Journal canadien de mathématiques, 11(??):397–408, ???? 1959. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Mahowald:1959:PMS

[570] Mark Mahowald. On a paper of Maurice Sion. Canadian Journal of Mathematics = Journal canadien de mathématiques, 11(??):409–415, ???? 1959. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Ellis:1959:GMT

[571] H. W. Ellis. On generalized Morse– Transue function spaces. Canadian Journal of Mathematics = Journal canadien de mathématiques, 11(??):416–426, ???? 1959. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Jenkins:1959:TP

[572] James A. Jenkins. On a type problem. Canadian Journal of Mathematics = Journal canadien de mathématiques, 11(??):427–431, ???? 1959. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Rooney:1959:SPF

[573] P. G. Rooney. On some properties of functions analytic in a half-plane. Canadian Journal of Mathematics = Journal canadien de mathématiques, 11(??): 432–439, ???? 1959. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Fulkerson:1959:NFF

[574] D. R. Fulkerson. A network-flow feasibility theorem and combinatorial applications. Canadian Journal of Mathematics = Journal canadien de mathématiques, 11(??):440–451, ???? 1959. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Sherk:1959:RMS

[575] F. A. Sherk. The regular maps on a surface of genus three. Canadian Journal of Mathematics = Journal canadien de mathématiques, 11(??):452–480, ???? 1959. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Amemiya:1959:CML

[576] Ichiro Amemiya and Israel Halperin. Complemented modular lattices. Canadian Journal of Mathematics = Journal canadien de mathématiques, 11(??): 481–520, ???? 1959. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Marcus:1959:TGS

[577] S. Marcus. Sur un théorème de G. Szekeres, concernant les fonctions monotones et convexes. (French) [On a theorem of G. Szekeres concerning monotone convex functions]. Canadian Journal of Mathematics = Journal canadien de mathématiques, 11(??):521–526, ???? 1959. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Brauer:1959:SRU

[578] Fred Brauer. Some results on uniqueness and successive approximations. Canadian Journal of Mathematics = Journal canadien de mathématiques, 11(??): 527–533, ???? 1959. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Wyman:1959:ABL

[579] Max Wyman. The asymptotic behaviour of the Laurent coefficients. Canadian Journal of Mathematics = Journal canadien de mathématiques, 11(??):

534–555, ???? 1959. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Michael:1959:CSC

[580] Ernest Michael. Convex structures and continuous selections. Canadian Journal of Mathematics = Journal canadien de mathématiques, 11(??):556–575, ???? 1959. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Nijenhuis:1959:NHR

[581] Albert Nijenhuis. A note on hyperconvexity in Riemannian manifolds. Canadian Journal of Mathematics = Journal canadien de mathématiques, 11(??): 576–582, ???? 1959. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

deBranges:1959:LOP

[582] Louis de Branges. The a-local operator problem. Canadian Journal of Mathematics = Journal canadien de mathématiques, 11(??):583–592, ???? 1959. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Kennedy:1959:THC

[583] P. B. Kennedy. On a theorem of Hayman concerning quasi-bounded functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 11(??):593–600, ???? 1959. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Hayes:1959:SPU

[584] Charles A. Hayes, Jr. Some principles underlying the construction of measures. Canadian Journal of Mathematics = Journal canadien de mathématiques, 11(??):601–613, ???? 1959. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Schuster:1959:PNP

[585] Seymour Schuster. Pencils of null polarities. Canadian Journal of Mathematics = Journal canadien de mathématiques, 11(??):614–620, ???? 1959. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Tutte:1959:SCG

[586] W. T. Tutte. On the symmetry of cubic graphs. Canadian Journal of Mathematics = Journal canadien de mathématiques, 11(??):621–624, ???? 1959. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Edge:1959:QTR

[587] W. L. Edge. Quadrics over GF(2) and their relevance for the cubic surface group. Canadian Journal of Mathematics = Journal canadien de mathématiques, 11(??):625–645, ???? 1959. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Coxeter:1959:PTR

[588] H. S. M. Coxeter. Polytopes over GF(2) and their relevance for the cubic surface group. Canadian Journal of Mathematics = Journal canadien de mathématiques, 11(??):646–650, ???? 1959. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Whittaker:1959:SHG

[589] James V. Whittaker. On the structure of half-groups. Canadian Journal of Mathematics = Journal canadien de mathématiques, 11(??):651–659, ????

1959. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Reiner:1959:CNR

[590] Irving Reiner. On the class number of representations of an order. Canadian Journal of Mathematics = Journal canadien de mathématiques, 11(??): 660–672, ???? 1959. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Goffman:1962:CPS

[591] Casper Goffman and G. M. Petersen.

Correction to the paper: “Submethods of regular matrix summability methods”. Canadian Journal of Mathematics = Journal canadien de mathématiques, 14(??):384, ???? 1962. CODEN CJ-MAAB. ISSN 0008-414X (print), 1496-4279 (electronic). See [340].